A Bibliography of Publications in Information Processing Letters: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 April 2024 Version 3.14

Title word cross-reference

#10015 [Day80a].

0 [RV79]. 1 [Gab76b, RV79]. 103 [Sch78]. 14 [Nos74]. 2 [Bro79c, Yao79b]. 21/2 [Kar73]. 3 [Bro79c, CM75]. 3NF [JF82]. 5 × 5 [Sch78]. 68 [LLW79]. 9 [Nos74]. ^{1,5} [Bud79]. B [Sam76]. c^{n} [Rob79]. d [SF79]. F [Müh79]. G [Rou73]. H [Rou73]. K [RW76, Roz76b, Yuv75, Zak82, Pet80b, SF79, Szw79]. L [Bui77, Bui79, Day80a, Ruo75]. LL(1) [FTM79]. max(m, n) [Rou73]. N [Pen75, Bha80, Gab76b, IMR80, Man76b, Urb80]. n^{2} [Cor71]. N^{2.81} [Yuv76a]. n × n [BCRL79]. ν [Cul78]. O(k^{2} log n) [Pet80b]. O(k^{3} log n) [Pet80b]. O(log n) [Sha79, Sho78, Urb80]. O(N) [HG77]. O(n^{2.7799}) [BCRL79].

O(n log n) [Röh82]. O(V^{3}) [MKM78]. O(|E|) log log |V | [Yao75a, Yao75b]. ω [Fey76]. Ω((n/ lg n)^{1/2}) [RV79]. Ω(n log n) [Avi80, van80]. P [DLR79, CL80]. QY [Mak79]. S [Day80b]. T [DKZ77]. θ(log log N) [Fra79b]. V^{2} [HT71]. Y^{T} Q [Mak79].

-ary [Zak82]. -connected [Gab76b]. -cube [Bha80]. -dimensional [Yuv75]. -Events [Fey76]. -factors [Gab76b, Müh79]. -Grammars [Pen75]. -isomorphism [Yao79b]. -iteration [RW76, Roz76b]. -Local [Cul78]. -matroid [CM75]. -modes [LLW79]. -polynomial [DKZ77]. -Reconstructability [Man76b]. -stability [Day80b]. -stable [Bui79, Day80a, Bui77]. -Trees [Sam76].

0L [Roz72a].

1 [Han74]. 10 [Avi80, Hel81, Llo80a, Mei81, vW80a]. 11 [Mor81]. 12 [Bin82, Bv82, Kri82]. 13 [NI80]. 14 [Kuč83, Tar86, YWL85].

56 [Day80a].

6 [Day80a]. 68 [Bak72, BCDL72, Mee76].

7 [Fou79, Gon79, Ov80, Yuv79].

8 [APT82]. 80b [APT82, Fou79, Ov80]. 80e [NI80]. 81a [CM80, Llo80a, vW80a]. 81k [Mor81]. 82d [Bv82]. 82i [Bin82]. 82j [Kri82]. 83g [Kuč83, YWL85].

9 [And80, CM80].

A. [Ov80]. Aanderaa [Ill78]. Aanderaa-Rosenberg [Ill78]. absence [CM79a]. Abstract [Arn74, FHMS79, Ber78]. acceptors [IT79a, NI80, SK77]. access [BN78a, BN78b, DD76b, GR76, HB77, Kuč82, Kuč83]. Accessing [Gun82]. Acquisition [Bra72]. adaptive [Jav77]. Addendum [AT79, Bro79b, Gra77b, Kuč83]. adder [Joh80]. Addition [RK75b, RK75a, CRT79, CRM75a, CRM75b, Gra80, Hug78]. additions [RV79]. additive [HV75, Krá72a, Krá72b]. Address [Fis74]. Address-Order [Fis74]. advancement [Jav77]. AFL [Boa73]. aggregates [FHMS79]. AL1 [Bra78]. algebra [Par78]. Algebraic [MI73, BM82, Boa73, CE78, DL78a, Dem82, Lau80, SS77, Win77, Win78]. algebras [BC79]. ALGOL [Bak72, BCDL72, LLW79, Mee76]. Algorithm [AKS78, And78, CM77, DR76, GS87, HG77, KJ78, MA80, RE79, Roh78, AT78b, AT78a, AT79, Akl79, And79, And80, AS78b, APT79, APT82, BY76, Bin82, CL80, Che78, CS77, CLW80b, Cla75, CGH75, CRT79, Cor71, Der80, Dew74, Dye80, FH79, FZ82, GT78, Gav75, GR77, Gra72, Han76a, Han76b, Han78, Han76c, HT71, Hue78, Ita79a, Jav77, Jon79, Jov74, KHTY78, KST82, Knu77, KP72, Kun76, LD80, Loi80, MKM78, MA79, MV74, NMM78, Ov80, PS80, PS81, Pet80b, Rei78, Rob73, RF73, Rou73, Sam78, ST79, Sch78, Shi79b, Ste75, Tar74a, Tar74c, Tar82b, Tsa81, Urb80, WS82, Yao75a, Yao75b, Yuv76a]. Algorithms [AS78a, Cof73, Coh75, Dwy73, Gal76, GIG77, GI79, NM82, AR80, Ber82, BMT79, Bła77, Fre80, HN79, Ko82, Kob74, MH79, Mit79, ST77, Sik72]. Aliasing [Sch79]. allocation [Gel71, Han76c, Kro73, Rob80]. almost [BY76]. Alphabetic [Wes76]. alphanumeric [SRK78]. Alternate [Red77]. alternating [RS77]. ambiguity [Tak74]. Ambiguous [Tar82a]. among [Sch79]. Analysis [BS75, CM77, HWH76, IT76, MA80, SF79, Tsu78, Vik79, Yao80, CFG72, Fre80, OO72, Tar82c, Tar86]. Analyzers [KŁ76]. anomaly [Lam76, LT75, Tri76]. APL [EPS82]. Applicable [Reu76]. Application [de 78, Akl78, HN79, Man76a, PD79, Sta71]. applications [Bar73, DD76b, DKP77]. Applying [Bof73]. approach [Cza78a, MN72, PL79]. approximate [BCRL79]. approximates [MZ79]. Approximating [MG73]. approximation [CS77]. approximative [Noz79]. April [Avi80]. Arc [BB79]. ARCADE [BKH78]. Architecture [LR80, BKH78]. Area [PV80, Bre80]. Area-Time [PV80]. areas [DD76b]. Arithmetic [Cho82, Bar76, PM75, Sha79, Sho78]. arrangement [ten80]. Arrangements [KS74a, KS74b]. Array [RM73, Wan76a, Wan76b]. ary [Zak82]. Ascending [BWF^{+}79]. assertional [Bra78]. assigning [Mee76]. assignment [AR79, AR80, Set73]. assisted [DDM^{+}75, LC76]. associated [Pău79]. Associative [Min73, Ore82, Tes79]. assure [FL82]. Asymptotic [Dob82, Gon78, Gon79]. attack [Hel81]. Attribute [GW78, RB82]. Author [Dob78a]. authorization [PP80]. Automata [ITN78, RM73, UMS82, Iba74, IT80, Jon77, Kin80, NI80, Rei79a, Tza72b]. Automatic [CFG72, CK77, KŁ76, Gin82]. Auxiliary [Dwy73, Rob73, Sik72]. Average [SF79, Chr72]. avoidance [Cel79]. avoiding [DL78b, KF71a, KF71b]. Axiomatic [Sou79].

B [Bv82]. backtrack [HN79]. backtracking [All78]. Balanced [SF79, Bay72, Bro78, Bro79b]. bandwidth [Rob75]. Base [BB76b, Gra77a, FG79]. based [Lei80]. Bases [van79b]. basic [Wha72]. basis [Hue78]. Batch [Lyo79, Wil80a]. batching [LM78, Wil80a]. be [Kei75]. behavior [Dob82]. behaviors [NMM78]. Behaviour [Tsu78, Win77, Win78]. between [ER79a, Gav77, MC80, Rot75, Ruo75, SS77, VL79]. bilinear [Ber82]. bilinearity [Pro76]. bin [Fre80]. Binary [Ada73, Lei79, MOS76, Min73, NPWY72, Van79a, Bre80, Cla75, Han76a, Han76b, HR76, KM79, Kor82, LS82, Mor79b, Rob73, Rot75]. binding [Pri79]. Bini [Tsa81]. Binomial [Ram77]. bipartite [Gab76a]. Bisection [RE79]. bit [Jak78, Jak82, Teu78]. bit-vector [Jak78, Jak82]. bit-vectors [Teu78]. Bits [Lin73]. BL [Bro80a]. Boas [Avi80]. Bodies [HM73]. Boolean [APT82, PS81, APT79, PS80]. bottom [IT79a]. bottom-up [IT79a]. Bound [Hir78, PS78, Avi80, Ban80, Bro80a, FL82, Kun73, Nie75b, Ple79, RV79, TW73, van80, vS74]. Boundary [HM73, MG73, Han76c, MN72]. boundary-tag [Han76c]. Bounded [Fis74, Wei75, Iba74, LR74, Vel79]. bounded-reversal [Iba74]. Bounds [CM75, Leu79, Llo80b, NPWY72, BC79, BMS80, Gab76b, Llo80a, Sel78, vW80a]. Boyce [KM80, Osb79]. Boyce-Codd [KM80, Osb79]. Bracket [Meh76]. Bracket-languages [Meh76]. branch [Bro72, TW73]. branch-and-bound [TW73]. branching [Shi79a, Tar74a]. bridges [Cor71, Tar74b]. Bubble [Dob80, CLW80a, CLW80b]. Buffer [AL75, CHCP74, Vik79, Dij72]. buffered [RV77]. buffering [Tza72a]. Bui [Day80a]. Burton [Dob78a]. Bykat [Ov80].

Cahit [BL78]. calculate [Sho78].

calculation [Ste75]. calculus [Mer78]. Can [FW79b]. Candidate [Dem78, YJ76]. canonical [CPR71, Hug78, Sam78]. capacity [Pău78]. Cardinality [Hos75, Kan80]. Carlo [Ber82]. cascade [Han76a, Han76b]. Case [SF79, Fra77, Röh82]. categorical [Kar71]. category [WS82]. catenative [Kob77]. Catenativity [Cul78]. Cellular [UMS82]. centroid [KA75]. Certain [GTY78, APT79, APT82, Bar73, CK77, Mis79, Rot75]. chain [MT75]. Characterization [Cza80, Kob77, Win77, Win78]. Characterize [DS78]. Characterizing [Vel79]. cheaply [Mor79b]. check [AT78a, AT79]. Checking [Gun82, Ita79a, Shi79b, Val76]. choosing [War73]. chordality [Gav75]. chromatic [Law76, RF73]. cigarette [Red77]. circuits [DZ71, ZD71]. Circular [BB79, FW79b, Boo80, Ita79a, Shi79b]. Circular-Arc [BB79]. clarification [VL79]. Class [CRM78, FS73b, MI73, AM77, CK77, CRM75a, CRM75b, Kit77]. classification [Sol72]. clique [Nie75b, Sel78]. cliques [BB76a]. closest [Dye80]. Closure [Cof73, Boa73, Han76a, Han76b, IT79b, Mun71]. cluster [OO72]. clustered [Teu78]. clustering [Tar82b]. Codd [KM80, Osb79].

Code [Han80, Ban77, WB79, Kay76]. codes [Kri82, MR81, Vel79]. Coding [HG77, Jak78, SRK78]. coefficient [Ste75]. coefficients [Ram77]. coercions [Ban77]. Colbourn [CM80]. Collecting [FW76a, FW77]. Collection [Fis74, Cam74, Hoa74]. collector [Zav75]. collectors [TG78]. colored [Pet80a]. coloring [LM78, MH79, RF73]. colouring [Wil80a]. combinatorial [CK77, PD79]. combined [ST79]. Comments [Akl78, AJV86, Avi80, Boc76, Bur78, Day80a, DS79, FK79, Fou79, Lam76, Wür77, Dob78a, Ov80]. Common [Hir78]. communicating [CM79a]. communication [Hel81, LM80b, Mei81]. communications [CRT79]. Commutation [DS78]. commutative [Mak79]. commutatively [Kri82, MR81]. Commutativity [Pro76]. compactifying [TG78]. compacting [Zav75]. compaction [Jon79]. Compactness [FS73b]. Comparison [Laz75, Bar73, Ruo75, Tsi71, War73, VL79]. comparisons [Nos74, War73]. Compilation [GW78]. Compile [Gun82, Han74]. Compiler [Arn74, KŁ76]. Compiler-Compiler [KŁ76]. compilers [Tur77]. complement [Ban80, Rob79]. Complete [Lev75, Wil77, ER79a, HR76, Lip77, NMM78, OO72, SG79, Van75, Yao79b]. completeness [Mat77b, Ple79]. completion [Kun76]. complex [vT72]. Complexity [BMS80, Cof73, GI79, JKS79, Pea75, Akl78, Ban80, BSW77, Bha80, BCRL79, BC79, CE78, Col79, CM80, DV76, HV75, IMR80, Joh80, JF82, Kan80, Kun73, Law76, Man76a, Noz79, Oul79, PR77, Röh82, Val76, YJ76, de 75]. Components [NM82, Arj76, Che78, Tar74c, Tar82b]. compression [Jak78, Jak82, Teu78]. Compromising [van79b]. Computation [Arn74, DM79, Ers77, Kuč82, Kuč83, Zuc77]. computational [Dei74, Kan80, Sch82].

Computations [Pea75, DS80]. compute [Mak79, RV79]. Computer [Her73, Sri74, Sri75, BKH78, DDM^{+}75, LC76, Ste75, Tza72a]. computer-assisted [DDM^{+}75]. Computer-Constructors [Her73]. computers [PD79]. Computing [Cof73, GL80, Val80, Abe79, AR80, IMR80, Pet80b, Urb80]. concerning [CKH79, Yao79a]. Concurrent [Bro80b, Chr76, HB77, Bes79, Han75]. Conditional [DL73, KM79, Sam78]. conditionals [Red77]. Conditions [MG73, Wil77]. configurations [Wil80b, vS74]. conflicts [Kuč82, Kuč83]. conjecture [Ill78, Yao79a]. Connected [NM82, Arj76, Che78, Gab76b, Yuv73]. conquer [BS78]. Consecutive [EL76, LP76, ten80, Lip77]. conservation [Hoa73]. consistency [FL82]. Consistent [MI73]. constants [GM79]. constrained [Gab76a]. Constraints [LF76, Bła79, Kan80, Leu79, Sło79]. Constructing [HR76]. Construction [Chr76, Ang79]. Constructors [Her73]. consumer [Wür77]. Containment [FG78]. content [CFG72]. Context [FG78, Sli82, AM77, BN78a, BN78b, BCDL72, CMM73, Jaf78, Lin76, Man76b, MR71, Roz72a, Roz76a, Sok78, Tak74, Ukk79, Weg75]. Context-Free [FG78, Sli82, AM77, BCDL72, Jaf78, Lin76, Man76b, Roz72a, Sok78, Tak74, Ukk79]. continued [Akr79]. continuous [Abe79]. contractable [SY72]. Control [DT75, CKH79, DD76b, DL78b, Kru79]. Convergence [MG73]. conversion [Coh73]. Convertibility [ter80]. Convex [And78, Byk78, GS87, Jar73, KJ78, AT78b, Akl79, And79, And80, Avi80, Bro79a, DS79, Dev80, FK79, Fou79, Gra72, LP78, MA79, Ov80, YL79, van80]. Copying [ES76, Cla75]. coroutine [Han76c, Pri76]. correcting [SRK78]. correction [Che78, CM80, Kra80]. correctness [Bra78, DY75, Sou79, Upf82]. correspondence [Rot75]. Corrigendum [And80, Tar86, Yuv79]. Cost [Mid74, Sło79]. Cost-minimal [Sło79]. Cost-Oriented [Mid74]. costs [JS80]. Counter [GW78, Boa73]. Counter-One-Pass [GW78]. counterexample [Ill78]. Counters [Wei75]. Counting [CRM78, FW79b, Jaf78, BMT79, Mat79]. Covering [BB79, Osb79, Gav77, KM80, Ukk79]. covers [Mou80]. CR [SBS79]. Creates [FW76b, FW79a]. Critical [KL78, KM79, Wha72]. Crossvalidation [Chr75]. cryptanalytic [Hel81]. Cryptography [CM79b]. cryptosystem [Hel81, LM80b, Mei81]. cube [Bha80]. Cubic [Wil80b]. Cycle [KL78, Loi80, Ple79]. cycles [Jov74]. cyclic [IT79b, RV77, ten80]. Cyclomatic [Oul79].

D [Bv82, Day80a]. D.P. [Dob79]. D0L [Cul78, LR74, Lin76, Ruo79]. D0L-ness [Lin76]. D0S [ER80]. Data [BB76b, Bra72, FW76b, FW79a, Gra77a, Loi79, SG77, Sol74, van79b, AG75, BR75, Ber78, Che79, FG79, Lao79, PL79, SRK78, Vár73, vT72]. Data-Bases [van79b]. Database [Gra79b, RB82, Luk79]. databases [Kan80, LD80, SR79, Yao79a, ter80]. dates [Sło79]. DB [FG79]. Deadline [Bła79]. deadlines [Bła77]. Deadlock [CM79a, Cza80, BN76, Cel79, JS80]. Decidability [Cul78, CRM75a, CRM75b, Cul79, Gra80, Ruo79]. decidable [Lin76]. Decision [Cza80, HR76, IT80]. Decomposable [Ben79, vW80b]. deduction [Far82]. defined [GL80]. Defining [RB82]. definition [Boc76, CG76, Mad77]. degree [Gab76a, Ple79]. degree-constrained [Gab76a]. Delaunay [MZ79]. delayed [Erk80]. demand [AS78b, BM77]. Denning [Bv82]. dependence [Kit75]. dependencies [GM82, YJ76]. dependency [Luk79]. dependent [Kei75]. depending [MS72]. depth [Hen79, McC78]. Derivation [Pet80b, Gon78, Gon79]. derivations [Jaf78]. derivative [Kun73, ST77]. describing [Sli82]. Description [Sol74]. Design [AL75, BB76b, Chr76, MA80, Tza72a]. Destroys [FW76b, FW79a]. detecting [vS74]. Detection [KL78, BN76, DS80, FTM79, JS80, ST79]. determination [Avi80, BES82, Mun71, Nie75b, van80]. Determining [And78, Tar72, Cor71, Gra72, HT71, Kit77, Rou73]. determinism [All78]. Deterministic [Coh75, FG78, UMS82, ER73, Jon77, Kin80, Roz72c, Roz72b, SK77, vS74]. development [CKH79]. developmental [Kob77]. deviations [KP72]. Device [Sol74, Han74]. diagnosis [Tur77]. Diagonals [MRK76]. diagrams [Bro79a]. Dictionaries [MOS76]. Difference [MG73, MS72, Urb80]. differences [HSB73]. differential [Bui77, Bui79, Day80a, Wha72]. differentiation [AG75, CE78, Vár73]. diffusing [DS80]. digraph [Arj76]. digraphs [Ple79]. Dijkstra [Knu77, NMM78, Vv72]. Dimensional [ITN78, Bir77, Fre80, IT80, Ita79b, NI80, Yuv75]. dimensions [And79, And80, Byk78, DS79, Fou79, Ov80]. Diophantine [Hue78]. Direct [Roz72a]. Directed [Kam75, Coh73, Mun71, Shi79a, Shi79c]. disciplines [Hoa73]. discrete [Jav77, Kaw80, KP72]. discrete-time [KP72]. discriminated [BS79]. Disjoint [BB79, Shi79a, Tar74a]. disk [CT73]. distinct [Szw79]. distributed [AR79, AR80, DL78b, GR76, Kru79]. Distributive [Dob78b, HK79, AN80, Bur78, Dob78a, JKS79, Nog87]. divergence [Kit75]. Divide [BS78]. Division [CM77, Bar76, Dem82, PM75]. division-free [Bar76, PM75]. do [Oul79]. Document [RBE^{+}72, Tes79]. Domination [NW78, CGH75]. double [Nog87]. down [CKH79]. DPA [vS74]. Drift [RK75b, RK75a]. Driven [FW76b, FW79a]. dual [BR75]. duality [Mak79]. due [Sło79]. dynamic [Gel71, Kro73]. Dynamics [RB82]. Dynamization [vW80b].

E0L [Gra77b, AM77, Gra77c]. Earley [MV74]. economic [CRT79]. ECT [SY72]. Edge [Shi79a, MH79, Tar74a]. edge-coloring [MH79]. Edge-disjoint [Shi79a, Tar74a]. editing [Sel77]. editor [Kri82]. education [BKH78]. effects [CG76]. Efficient [And78, Dob80, KJ78, Mun71, Pap79, vN82, And79, And80, BMT79, Ghe75, Gra72, ST79, War73]. eigenvalues [BES82]. Elastic [HM73, Rem75]. Elasticity [MG73, MS72]. electronic [PD79]. element [Urb80]. elementary [Ste75]. Elements [van79b, Mer77, Nos74, Rei78]. embedded [Ros79]. Emde [Avi80]. emptiness [ER80, Rob79]. Enclosures [DD76b]. entropy [Chr75, Sek75]. Environment [Fis74, CFG72, Ita79b, LC76]. Environments [FW79b, Han76c]. equality [Gav77, Sch71]. equation [Urb80, Wha72]. Equations [HM73, MI73, dd75, dd76a, Akr79, Bui77, Bui79, Day80a, Hue78, MS72, PS80, PS81, Ruo78, WS82]. Equilibrium [HM73, Gel71]. Equivalence [FG78, Bay72, BMS80, EO82, GM82, Hug78, Ita79a, Kob74, Roz72a, Roz72c, Roz72b, Ruo79, Sam78, Shi79b]. equivalence-checking [Shi79b]. equivalent [Kri82, MR81]. Errata [Llo80a, vW80a]. Erratum [APT82, BN78a, Bui79, CRM75b, FW77, FW79a, Gon79, Han76b, Har79a, Hor72a, KF71a, KS74a, Krá72a, Kuč83, PS81, RK75a, Roz72c, Sri75, Wan76a, Win78, YWL85]. Error [Kra80, Tur77, Wür77, CS77, FTM79, Ghe75, SRK78]. error-correcting [SRK78]. estimates [Dob82]. estimation [Ber82]. estimator [Yuv77]. ET0L [Pen75, Roz76a, Van75]. ET0L-Grammars [Pen75]. ET0L-languages [Van75]. ETOL [YWL85, YWL82]. EUCLID [Sch79]. Euclidean [BM79]. Evading [RK75b, RK75a]. evaluate [PM75]. Evaluating [BM71, Val76]. Evaluation [Jak82, Bar76, CS77, Chr72, Han80, Kun73, Noz79, ST77, Tou72]. evaluations [PD79]. evaluator [RS77]. Even [SG79, Ebe80]. Events [Fey76]. example [Kar73]. Examples [MTT73]. exchange [Cah77]. exclusion [AJV86, Mor79a]. Execution [GL75, LF76]. Exhibit [Tsu78]. Existence [Osb79, CPR71]. exit [CKH79]. expansion [Zav76]. Expected [DR76, BS78, Dob82, Fra79b, JS80, NMM78]. experiment [KK72, Mad77]. experimental [AG75, Mat77a, Vár73]. Experiments [Sal73]. Exploiting [AS78a]. Expression [Chr72]. Expressions [Cza78b, Bar76, Gon78, Gon79, GM82, Kar71, PM75, Rob79]. expressive [Mer78, Par78]. extended [Rob79]. Extendible [Yao80]. extending [Kob74]. extensible [Sol72, SY72]. extensible-contractable [SY72]. extension [FZ82, MZ82, Vv72]. Extensions [Cam78]. extremely [Krá72a, Krá72b].

F [Bv82, Mor81]. F0L [Ruo79]. F0L-D0L [Ruo79]. facilities [Han74, Mad77]. Factoring [Sha79]. factors [Ebe80, Gab76b, Müh79]. Fairness [Cza80]. families [Szw79]. family [AM77, Boa73, Kri82, MR81, PP76, ST77]. fan [Bar76]. fan-in [Bar76]. Fast [AKS78, BD80, Gal76, MAP76, PS78, SG76, Sik72, de 78, AT78b, Ber82, CS77, Cla75, Dye80, Hor72a, Hor72b, Jon79, Shi79b, Zav75]. Feature [Tou72, Kit75]. Features [GW78]. feedback [GT78, SG79, Tza72a]. Few [van79b]. Fibonacci [GL80, Pet80b, Sho78].

Field [Gun82]. fields [Cho82]. FIFO [ES77]. FIFO-optimal [ES77]. file [BR75, HB77, SBS79]. files [Sta71]. filter [KP72]. Find [DR76, Lao79]. Finding [ER79a, HG77, Joh75, Yuv75, Yuv76b, Arj76, BD80, BSW77, Che79, Dev80, GT78, Jov74, KST82, Loi80, MKM78, MA79, Tar74c, Tar74b, Yao75a, Yao75b, YJ76, Yuv76a]. Finite [And78, CHCP74, ITN78, Jar73, Byk78, Cho82, DS79, Dij72, Fou79, Gra72, HSB73, IT80, LLW79, NI80, Ov80, Pău78, Rei79a]. First [GL75, RV77]. first-in-first-out [RV77]. Fixed [LNS82, BSW77]. fixed-radius [BSW77]. Floating [RK75b, RK75a]. Floating-Point [RK75b, RK75a]. floor [Sch82]. flows [MKM78]. folk [LNS82]. FOR-loops [FHMS79]. forest [van77]. Form [Osb79, Nij79, Sam78, YWL82, YWL85]. Formal [MTT73, Upf82, Kaw80, LLW79, Mad77, Tal77]. Formalization [GW78]. forms [Ber82, CPR71, KM80, Roz72a, Sam78]. formulae [McC78, SK77]. formulas [APT79, APT82]. FORTRAN [DDM^{+}75]. forward [Fra77]. fractions [Akr79]. Free [FG78, Sol74, AM77, AJV86, Bar76, BCDL72, Che79, EPS82, Jaf78, KA75, Lin76, Man76b, Mor79a, PM75, Roz72a, Sli82, Sok78, Tak74, Ukk79]. Full [San76]. Fully [Kas73]. function [MZ82, Moo79, Sch82, Sek75]. Functional [Łuk77, YJ76]. functionals [SS79]. Functions [San76, Abe79, Eng78, GL80, Hen79, KR76, Shy76]. fundamental [Jov74]. Further [Ov80]. future [Chr75].

G. [Akl78]. game [Lou80]. Garbage [Fis74, FW76a, FW77, Cam74, Hoa74, Jon79, TG78, Zav75]. General [Mer71, DL78b, Hoa73]. Generalization [Lei79, HSB73, Knu77]. Generalizations [DKP77]. Generalized [Bud79, Dem82, MI73, Reu76, Ill78, Müh79]. Generate [KS74a, KS74b, Hue78]. Generating [CRM78, MAP76, CL80, Ist78, Sek75]. Generation [KŁ76, Ple75, RBE^{+}72, Sri74, Sri75, Zak82]. generative [Pău78]. generator [Krá72a, Krá72b, WB79]. Genetic [TG78]. geometrical [Akl78, Man76a]. geometry [PD79]. Geschke [Mat77a]. GF [CL80]. Ginsburg [Mau71]. Given [BWF^{+}79, YJ76]. global [Mat77a]. go [KF71a, KF71b]. good [ST77, Tar74a]. Graceful [CC75, BL78, Cah77]. Graham [GS87]. Grammar [CRM78, Pău79]. Grammars [GW78, How78, MTT73, Pen75, RB82, Tar82a, Wan76a, Wan76b, CMM73, ER79b, Ghe75, Jan79, JK75, Kra80, Łuk77, Man76b, MR71, Nij79, Pău78, Roz72a, Roz76a, RW76, Roz76b, Sli82, Sud77, Sud78, Tak74, Ukk79]. grammatical [Mou80]. Grant [PP80]. Graph [Yao79b, Bha80, Cor71, Mat79, Mun71, Nie75b, RF73, Rou73, Sel78, Tar74b]. Graphics [RBE^{+}72]. Graphs [Bof73, BB79, Kam75, Kas73, Sri74, Sri75, DKP77, Ebe80, Gab76a, Gab76b, Gav75, Han74, Hen73, HT71, LM78, MH79, Mit79, Müh79, Nie75a, NMM78]. Gray [Kay76]. Greedy [MZ79]. Greibach [Nij79]. group [Bv82, DS81, Tar72]. groupoid [Tar72]. Growing [Kor82]. growth [Kar73].

Hamiltonian [Kun76, Ple79]. handling [Ghe75, Sta71]. Hanoi [BL80]. hard [DLR79, Ko82, Rob80, SK77, Sli82]. Hardcopy [RBE^{+}72]. harder [Mor81, Rom80]. harmonic [Zav76]. Hash [HG77]. Hashing [AKS78, GIG77, SG76, Yao80]. hazards [ZD71]. Heap [FW76a, FW77]. Height [MOS76, Wes76, Bro78, Bro79b]. height-balanced [Bro78, Bro79b].

Heuristic [AR80, Fin72]. Heuristics [BB76b]. hierarchical [Jak82, Tar82b]. hierarchies [Tsi71]. Hierarchy [Wan76a, Wan76b]. Higher [Coh75, Dob82]. Higher-Level [Coh75]. homogeneous [Hue78]. homomorphic [Nie75a]. homomorphism [ER79a]. Huffman [Jak78]. Hull [And78, GS87, Jar73, KJ78, AT78b, Akl79, Avi80, Byk78, DS79, Fou79, Gra72, MA79, Ov80, van80]. hulls [And79, And80, Bro79a, Dev80]. Hybrid [MA80, Röh82].

Identification [Jar73, JD74, SBS79, Tza72b]. images [Nie75a]. Immediate [FTM79]. Implementation [Cza78a, Cza78b, GL75, Boc76, GM79, Pri79]. implementations [KM79, Lau80]. Implementing [MOS76, HN79, Mis79, Set73]. Implicit [DM79]. Imply [Gal76]. Improved [PS78, RE79, AT78a, AT79, Bro80a, Gab76b, Rob73, vS74]. improvement [CL80, Han80]. Improvements [MAP76]. inbreeding [Ste75]. Includes [FW76a, FW77]. inclusion [Boa73, Kin80]. Independent [LF76, Sło79]. independently [ES77]. index [GR76, KHTY78, Pău78, Ram77]. indexed [DPS76]. induction [Gri72]. Inexact [Pea75]. infinite [Eng72, Yuv76a]. infinite-precision [Yuv76a]. Inform [And80, APT82, Avi80, Bin82, Bv82, CM80, Day80a, Fou79, Gon79, Hel81, Kri82, Kuč83, Llo80a, Mei81, Mor81, NI80, Ov80, Tar86, YWL85, Yuv79, vW80a]. Information [CHCP74, CM79b, Dij72, Hir78, Sal73, Sal71, Tal77, Tou72]. Information-Theoretic [Hir78]. initial [SG79]. Input [Tza72b]. Input-output [Tza72b]. Insertion [PE78, Dob82]. Inspected [Vik79]. instability [Bin82, Tsa81]. instruction [DDM^{+}75]. instructions [Bro72, Set73]. integer [RV79]. Integral [dd75, dd76a, Ram77]. integrating [MN72]. interaction [Ruo75]. interactive [FL82, Sal71]. interleaved [Rob75]. intermediate [Kun79]. intermediate-value [Kun79]. Internal [Erk80, Day80b]. interpolation [GR77, Hor72a, Hor72b, PR77]. interpolation-sequential [GR77]. Interpretation [Coh75, FW76b, FW79a]. intersection [CM75, EMK82, ER80]. Interval [Fra79a, IT76]. inverse [Kun77]. Inversion [PS78]. investigation [Mat77a]. involving [Ruo78, Zav76]. irreducible [CL80, Ist78]. isomorphism [BB76a, BMS80, HT71, Mat79, Yao79b]. iteration [RW76, Roz76b]. Iterative [MS72, MI73, Gin82, Sho78].

jobs [Sło79]. join [LD80]. Joshi [Kit75]. July [Hel81, Mei81]. June [Gon79].

Kam [Yao79a]. Keys

[Dem78, Bv82, DS81, YJ76]. Known

[van79b].

Language [Chr76, Bra78, Coy76, Dem82, Gra79a, Mer77, Ruo78]. Languages [Coh75, CRM78, FG78, Wan76a, Wan76b, Von75, AM77, BS79, Boa73, DPS76, ER73, ER80, Gra77b, Gra77c, LR74, Lin76, Meh76, Pău79, Pri79, Ruo75, Sok78, Sol72, Van75, Vel79, Yeh79, YWL82, YWL85]. largest [Boa73]. Latin [BMS80]. Lattices [Hos75, Reu76]. law [Hoa73]. Layer [Chr76]. layouts [Bre80]. lazy [Han80]. least [Boo80]. left [Nij79]. left-regular [Nij79]. leftmost [CMM73]. lemma [BM82, CPR71, Yeh79]. Length [BWF^{+}79, LF76, NPWY72, Krá72a, Krá72b, LR74]. Lenstra [Paj80]. less [Gra77b, Gra77c, van77]. Lett [And80, APT82, Avi80, Bin82, Bv82, CM80, Day80a, Fou79, Gon79, Hel81, Kri82, Kuč83, Llo80a, Mei81, Mor81, Ov80, Tar86, YWL85, Yuv79, vW80a]. Letter [Kri82]. Level [Coh75]. levels [PP80]. Lexical [KŁ76]. Lexicographically [Boo80]. limit [ER73]. Lindenmayer [How78]. Line [RBE^{+}72, CFG72, DKP77, NI80, Rou73]. Linear [DLR79, DR76, MH79, Mit79, MI73, APT79, APT82, BS78, BC79, Boa73, CGH75, FS73a, GT78, Hue78, Jon77, KHTY78, Kun76, Mat77b, MA79, MZ82, Rei78, Roz72a, Shi79c, Shy76, Tza72b, ten80, van77]. linear-time [APT79, APT82, GT78]. Lingual [Sal73]. link [BR75, CRT79]. LISP [CG78, FG79, de 78]. LISP-type [FG79]. List [Fis74, Lin73, Llo80a, Llo80b, vW80a, CG78, Leu79]. Lists [Fra79a, Lyo79, Fra79b, Han78, Ita79a, Shi79b]. Literate [Knu92]. LL [Ghe75, JK75]. LL-regular [JK75]. Local [Cul78]. Locality [IT76, Tsu78]. Locality-Transition [Tsu78]. locally [Kob77]. Location [LY79]. log [DLR79, GL80]. log-space [DLR79]. logarithmic [Meh76, van77]. logic [Far82, Har79a, Har79b, Hen79, Lei80, Mer78]. logical [Coy76, Han76a, Han76b]. Long [Fin72]. Long-Term [Fin72]. Longest [Hir78]. Loop [DY75, Che79]. looping [vS74]. loops [FHMS79]. lossless [LD80]. Lower [BC79, Hir78, Avi80, Ban80, Bro80a, FL82, Ram77, RV79, van80]. LR [Tar82a].

Machine [Bud79, Bro72, Dei74]. machines [SS77]. Manacher [Akl78]. Manage [FW79b]. management [Mer71]. many [BM71]. map [Wil80b]. Maps [Bud79, Wil80a]. March [Tar86]. Margulis [Ang79]. marked [Hen73]. marking [Der80]. Master [Bv82, DS81]. Match [BB76b]. Matching [Gal76, Akl78, BB76a, Bir77, Gav77, Man76a, Müh79]. Mathematical [Laz75, Loi79]. mathematics [CPR71]. Matrices [PV80, BES82, Col79, CM80, IMR80, Sch78]. Matrix [MRK76, PS78, BCRL79, CM80, Mak79, Mor81, Pău78, Pău79, Pet80b, Rom80]. matroid [CM75]. Mauceri [Kri82]. Max [Cam78]. Maximal [Wes76, Avi80, BB76a, Dev80, Mit79, van80]. Maximum [BWF^{+}79, Gav77, MKM78, McC78]. meaning [Coy76]. measure [Kit75, Oul79]. measures [VL79]. mechanical [Moo79]. mechanism [DD76b, DL78b, HJ71, PP80]. Median [Nos74, Ebe79]. meet [Bła77]. membership [Luk79, Van75]. Memories [DT75, Min73, CLW80b]. Memory [Fin72, Fra79a, SG77, CLW80a, Kuč82, Kuč83, RV77, Rob75, Sta71]. Merge [PE78, Erk80, FS73a]. Merging [Van79a, Bar73]. Method [SG76, Bar73, Bui77, Bui79, Day80a, Far82, GR76, HSB73, Hor72a, Hor72b, Kit77, LM78, Mat77a, Pet80b, Sok78, SS79, Teu78, Wil80a]. Methodical [BWF^{+}79]. Methods [MAP76, MI73, dd75, Day80b, MS72]. micro [JD74]. micro-operations [JD74]. Microprogramming [LR80]. Miller [Len79]. Min [Cam78]. Min-Max [Cam78]. Minimal [LS82, NM82, War73, Sło79]. Minimax [Bof73]. Minimization [DT75, SS79]. Minimizing [Chr75]. Minimum [BB79, CLW80a, Joh75, Roh78, Chr72, Dew74, Gav77, Tar82c, Tar86, Yao75a, Yao75b]. Mixed [MG73, Pri79]. modal [Far82]. Model [Arn74, Gra79b, Laz75, Cam74]. modeling [Tza72b]. Models [Reu76]. modes [LLW79, Sch71]. modification [Jov74]. modified [TW73]. module [AR79]. moments [Dob82]. monitor [Boc76]. monotone [Hen79, McC78]. Monotonic [Wan76a, Wan76b]. Monte [Ber82]. morphisms [Ruo78, WS82]. Morris [AJV86]. most [CRT79, War73]. MR [APT82, Bin82, Bv82, CM80, Fou79, Kri82, Kuč83, Llo80a, Mor81, NI80, Ov80, YWL85, vW80a, Day80a]. Multi [Sal73].

Multi-Lingual [Sal73]. Multidimensional [Ore82, RY77]. multigraphs [Shi79a]. multihead [Iba74]. Multiple [BN78a, BN78b, Sam76, Hel81, LY79, LM80b, Mee76, Mei81, Pri76]. Multiplication [MRK76, BCRL79, Mor81, Pet80b, Rom80, de 75]. multiplications [Sch78, Yuv76a]. Multiplying [PV80, Sch78]. multiprocessor [Bła77]. Multiprogramming [Laz75]. multisection [BES82]. multiserver [Tza72a]. multitape [Kin80]. multivalued [Hen79, Luk79]. multivariate [Val80]. mutational [Kra80]. Mutual [FW79b, AJV86, Mor79a, Tou72].

N [Bin82]. near [BSW77, BM79, Yuv75]. nearest [FH79, FK79, LP78, YL79, Yuv76b]. nearest-neighbor [FH79, FK79, LP78, YL79]. nearly [Kor82]. necessary [RV79]. neighbor [BM79, FH79, FK79, LP78, YL79]. neighbors [BSW77]. neighbours [Yuv75, Yuv76b]. Neither [MZ79]. ness [Lin76]. nestling [Kun77]. net [Gra79a, MS72]. Nets [DL73, Pet80a]. network [Cah77, CRT79, Shi79c]. Networks [Fin72, KL78, PV80, Tsu78, CM79a, MKM78]. No [AKS78, And80, APT82, Bin82, Bv82, CM80, Day80a, Fou79, Kri82, Kuč83, Llo80a, Mor81, NI80, Ov80, YWL85, Yuv79, vW80a]. Node [BB79, Gav77]. Noiselike [Fey76]. Non [Coh75, CRM78, FS73b, MI73, All78, MN72, Pro76, Ram77, SK77, Sok78, Win77, Win78]. non-commutativity [Pro76]. Non-Compactness [FS73b]. Non-Consistent [MI73]. Non-Counting [CRM78]. non-determinism [All78]. Non-Deterministic [Coh75, SK77]. non-integral [Ram77]. non-sequential [Win77, Win78]. non-unique [MN72]. noncommutative [Sch78]. nondeterministic [MC80, Upf82]. nonexistence [Ukk79]. Nonisomorphic [DKZ77]. nonlinear [Wha72]. nor [MZ79]. Normal [MAP76, Osb79, KM80, Nij79, YWL82, YWL85]. normalization [SR79]. Notational [Sol74]. Note [Fin72, GS87, ITN78, Jov74, Yao80, Abe79, Ang79, BM79, Bes79, Bv82, CE78, CKH79, CG78, CT73, CMM73, CG76, Der79, Dev80, Ebe79, Ebe80, Eng72, FH79, Gab76a, Iba74, IMR80, IT79a, IT79b, IT80, Jan79, Jon77, Kob74, KD73, KR76, Law76, Lei80, LM80a, Lou80, Mat79, Mei81, MR71, Mor81, Mou80, Noz79, Pet80a, RY77, Rei79a, Ros79, Row75, Row77, RW76, Ruo78, Set73, Shy76, Sud77, Sud78, Tak74, Tar74b, Tsi71, YL79, Yao79a, Yeh79]. Notes [Gon78, Gon79, KF71a, KF71b]. NP [ER79a, HR76, Ko82, Ple79, Rob80, SG79, Yao79b]. NP-complete [ER79a, HR76, SG79, Yao79b]. NP-completeness [Ple79]. NP-hard [Ko82, Rob80]. Null [Gra77a]. Number [Dem78, Bac82, CLW80a, CGH75, Gab76b, Krá72a, Krá72b, KD73, Kun76, Law76, Nie75b, RV79, RF73, Sel78]. Numbering [CC75]. numbers [GL80, Oul79, Pet80b, Sha79, Sho78, Zav76]. Numerical [HM73, Vár73, dd75, dd76a, AG75, Ber82, Bin82, Tsa81].

observation [ER79b]. observations [Bro79c, Ko82, Nie75b]. Obtaining [MI73, RF73]. occurrences [ST79]. odd [Ebe80]. off [Mis79]. omission [Bak72]. on-line [CFG72, NI80]. One [GW78, Lip77, MI73, UMS82, Boa73, Fre80, PD79, Tur77]. one- [Fre80]. one-counter [Boa73]. One-Pass [GW78]. One-Way [UMS82]. only [Sik72]. open [Sło79]. operating [Mad77, Tsi72, Var72, Zel74]. Operations [SG76, Sol74, Fra79b, IT79b, JD74, Mis79, Sho78]. operator [Lev75, Sud77, Sud78, Wil77]. operators [EPS82]. Optimal [AS78a, Cam74, MT75, PV80, Wes76, AS78b, BY76, ES77, HR76, KST82, KK72, Kor82, KP72, MZ79, RF73, Row75, Row77, ten80]. Optimisation [Mid74]. optimised [Ban77]. Optimization [Hoa74, Loz76, Mat77a, RS77]. optimizing [Mat77b]. Optimum [NW78, Kun79]. Order [Fis74, IMR80, Pet80b, TG78, van77]. order- [Pet80b]. ordered [Han78]. orderings [Der79, Shi79c]. organisation [SBS79]. organization [BR75, Zel74]. Oriented [Bay72, Mid74, SG77]. Other [CHCP74, Cul78, DD76b]. outerplanar [Mit79]. Outline [Bro72]. Output [FW76b, FW79a, Tza72b].

packing [Fre80]. Padded [Fra79b]. page [BM77, Ita79b]. page-on-demand [BM77]. pages [ES77]. Pagination [IT76]. Paging [AS78a, Ita79b, vT72]. pair [Dye80, Mak79]. Paired [Fra79a]. Palindrome [Gal76]. Palindrome-Recognition [Gal76]. paper [Bin82, CM80]. Parallel [BES82, Cza78b, Cza80, GIG77, Kuč82, Kuč83, MRK76, NM82, PS78, RM73, Bar76, Cza78a, Cza79, DD76b, Dye80, IMR80, JD74, Kaw80, Nog87, Set73, Val80, Zuc77]. Parallel/Sequential [RM73]. parameter [Wha72]. parameters [Tza72a]. parsers [FTM79]. Parsing [Tar82a]. Part [GL75]. Partial [BB76b, Gra79b, Eng78, Ers77]. Partial-Match [BB76b]. Partitioning [Dob78b, HK79, Van79a, AN80, Bur78, Dob78a, JKS79, Nog87]. partitions [Kay76, MT75]. Pascal [Han75]. Pass [GW78, Tur77]. Path [BB79, Cza78b, KL78, NPWY72, Mor81, NMM78, Rom80, Tar82c, Tar86]. paths [Ros79, Yuv76a]. pattern [Akl78, Bir77, DPS76, Man76a]. Patterson [FZ82]. PD2L [Kar73]. PD2L-system [Kar73]. pebble [Lou80]. performance [CT73, Sal71]. periodic [LM80a]. Periodically [Vik79]. Periodicity [HG77]. permutation [Cah77, CLW80a, CLW80b, Kun77, Rot75, SS79]. Permutations [Ple75, BL78]. persistence [Gra80]. Petri [DL73, Gra79a, Pet80a]. Phrase [All78]. PL [Han74]. PL/1 [Han74]. placement [ES77]. Planar [And78, Kam75, Gra72, HT71, LY79, LM78, MZ79, Ple79, Wil80a]. Plane [Jar73, MG73, BM79]. Point [RK75b, RK75a, LNS82, MZ79]. Pointer [Sol74]. Pointer-Free [Sol74]. pointers [Sch79]. Points [Jar73, BM71, Byk78, DS79, Fou79, LY79, Ov80]. policies [CT73]. Polygon [GJPT78, AT78a, AT79, Byk79, MA79]. Polygonal [EMK82]. polygons [FK79, LP78, YL79]. Polynomial [San76, Wei75, Akr79, CS77, DKZ77, Lip77, ST77, Sli82]. polynomial-time [Sli82]. polynomially [Van75]. polynomials [BM71, CL80, HV75, Noz79, RV79, Val80]. Pope [CM77]. ‘Possible’ [Luk79]. Post [Hug78]. Potential [dd75, dd76a]. power [Dei74, Par78, Sch82, Vel79]. power-bounded [Vel79]. pp [Avi80, Gon79, Hel81, Mei81, Tar86]. Practical [Dob79, HK79, DY75]. precedence [KR76, Lev75, Shy76, Sud77, Sud78, Wil77, Zim72]. precision [Yuv76a]. preconditioning [Hor72a, Hor72b]. predecessor [Han78]. predicate [Fra77, MC80, Mer78]. preemptive [LM80a, Row77, Sło79]. prefix [Kri82, MR81, Vel79]. prepaging [AS78b]. Preserving [Fis74, van77]. primality [Len79]. Primitives [DS78, DM79, Vv72]. Principle [dd76a, Ers77]. priori} [KP72]. Priority [Joh75, Var72]. Probabilistic [Fre80, Roh78, CL80, DL78a, Ko82, Lei80]. Problem [BWF^{+}79, Cam78, GS76, HM73, Hir78, MG73, Akl78, AJV86, Ban80, BN78a, BN78b, BL80, Dye80, ER80, FK79, Kin80, Kro73, LLW79, Lao79, Law76, LP78, Lip77, Loz76, Man76a, Mat79, Mau71, MV74, Mor81, Mor79a, Müh79, PD79, Ple79, RY77, Red77, Rob75, Rob79, Rom80, Roz72a, Roz72c, Roz72b, Ruo79, Sel77, SG79, YL79, Yuv73]. Problems [Ben79, Cza80, FG78, NM82, UMS82, dd75, dd76a, BMT79, BN76, CM75, CK77, Cza79, EO82, GTY78, Gra79a, IT80, Ko82, MN72, Sli82, vW80b]. procedural [Lau80]. procedure [And72]. Process [And80, APT82, Avi80, Bin82, Bv82, CM80, Day80a, Fou79, Gon79, Hel81, Kri82, Kuč83, Llo80a, Mei81, Mor81, Ov80, Tar86, YWL85, Yuv79, vW80a, AR80, Har79a, Har79b]. processed [Tal77]. Processes [Tsu78, CM79a, Var72]. Processing [Fis74, Loi79, Han74, Tza72a]. Processor [MRK76, PS78, AR79, Row75, Row77]. processors [Tes79]. producer [Wür77]. Producing [Ban77]. Production [Arn74, Ebe79]. productions [Jaf78]. products [DV76, Mak79]. Program [Chr76, FS73b, GL75, HWH76, IT76, KS74a, KS74b, Mid74, Wei75, Bes79, CKH79, DL78a, DY75, Mat77a, PL79, Sho78, Upf82]. Programmed [DZ71, ZD71]. Programming [Gri72, Knu92, DD76b, DLR79, LC76, Mer77, Rem75, Sok78, Sol72]. Programs [Bro80b, FW76b, FW79a, Che79, Coy76, Gin82, GTY78, Kaw80, LLW79, MC80, Mat77b, Mer71, Oul79, Sou79, SS77]. projections [GG82]. proof [Bes79, Moo79, Pri76, Yuv78, Yuv79]. proofs [CM79a, Roz72a, Sou79, Upf82]. proper [Boa73]. Properties [Cul78, Kas73, BN76, KA75, NI80]. property [LD80, SBS79]. Protection [Tsi72]. Proving [Bra78, DY75, Sam78, Sok78]. pseudorandom [Krá72a, Krá72b]. pumping [Yeh79]. pure [And72]. pushdown [Cul79, Iba74, Jon77]. pyramid [IT79a].

QT [Mer78]. quad [BS75]. quadratic [PS80, PS81, Tou72]. quantified [APT79, APT82]. quaternion [de 75]. quatric [Gra77b, Gra77c]. question [Van75]. Queueing [AL75, Hoa73]. Queues [Joh75]. Queuing [Tsu78]. Quicksort [Röh82]. quindiagonal [BES82].

Rabin [CL80, FH79]. race [ZD71]. radius [BSW77]. radix [Kei75]. radix-dependent [Kei75]. Random [MAP76, And72, BR75, Bro79c, CMM73, Roz76a]. randomized [Ita79a]. range [BS75]. rank [IMR80]. ranking [Zak82]. Rational [Rei79b, vN82]. rationale [AG75]. rationals [Pap79]. Reachability [Kam75]. read [Sik72]. read-only [Sik72]. reader [BN76]. readers [BN78a, BN78b]. ready [Bła79]. Real [UMS82, LM80a]. Real-Time [UMS82, LM80a]. Realization [Cah77, DL73]. really [Kar71]. Recognition [Gal76, Gra77b, Gra77c, vN82]. recognizable [Meh76]. recognize [Mit79]. recognizer [KD73]. recognizing [JF82]. Reconstructability [Man76b]. record [BS79]. recovery [BN76, Tur77]. rectangle [EO82]. Recursion [FW79b, May72, PP76]. Recursive [FW76b, FW79a, GI79]. Recursiveness [Wan76a, Wan76b, GMS75]. reduced [KP72]. Reducible [Kas73]. reducing [Bar73, Kar71]. Reduction [Ita79b]. Redundancy [EL76, ten80]. Reevaluation [And78]. referees [Ano78]. Reference [FW79b, KHTY78]. referenced [ES77]. referencing [HJ71]. refinement [And72]. Region [SF79, KM79]. Register [AL75]. regular [Cul79, JK75, Nij79, Rob79, Vel79, Yeh79]. Related [UMS82]. relation [JF82, MC80, SS77]. Relational [Gra77a, BB76a, Dem82, GM82, Kan80, LD80, Luk79, MC80, Par78, SR79]. Relations [DS78, Mer77, Bay72, Han76a, Han76b, LS82, vN82]. Relative [Val76, CT73]. release [Sło79]. Remark [Dei74, DL78a, NI80]. remarks [Akl79]. Remer [KD73]. removal [PP76]. Reply [Bin82, Dob78a, Mor81]. Representation [Ada73, GS76, Kam75, Min73, Sch71, DPS76, MV74, Rei79a]. representations [LS82]. representatives [Szw79]. required [PM75]. Requirement [EL76]. Requirements [Bra72, SG77, ten80]. requires [Rob79]. research [BKH78]. resolution [MZ82]. Resource [SG77, Bła79, Leu79, Sło79]. Resources [Llo80b, Llo80a, vW80a]. respect [SK77]. Restivo [Kri82]. Restricted [Wes76, CKH79, CMM73, Leu79, YWL82, YWL85]. result [Ban80, BL78, Yuv78, Yuv79]. Results [CM79b, Laz75, Cul79, Har79a, Har79b, Ruo75]. resynchronisation [Wür77]. Retrieval [BB76b, CM79b, EL76, LP76, Reu76, Sal73, Weg75, Lip77, Sal71, Tes79, ten80]. reversal [Iba74]. Revisited [Dob79, Der80]. right [Vel79]. ring [Mak79, RV79]. role [Sta71]. Romani [Mor81, Mor81]. Rosenberg [Ill78]. Rosenbrock [Day80b]. rounding [Kei75]. Rule [Bof73, Gel71, Kei75, Pri76]. rules [PP76]. run [HJ71]. run-time [HJ71].

S [Kri82]. Safe [Lau80]. safety [Cel79]. same [DKZ77]. Sardinas [FZ82]. saving [Mee76]. Scan [RBE^{+}72]. Scanning [Lin73, CT73]. Scatter [FW76a, FW77]. scattered [ER79b, MR71]. Scheduling [LF76, Llo80b, Lyo79, Bła77, Bła79, Leu79, LM80a, Llo80a, Row75, Row77, Sło79, Var72, vW80a]. Schemas [Cza80, FS73b]. Schemata [Wei75]. scheme [Bro78, Bro79b, SRK78]. Schemes [MG73, GTY78, JF82]. Schneider [Bv82]. Schorr [Der80]. Sci [NI80]. science [VL79]. scoping [Pri79]. Search [Fin72, GS76, Lei79, San76, SF79, Wes76, And72, GR77, Kor82, Pap79, PR77, Rei79b]. searches [BS75]. Searching [Ben79, Ore82, BD80, BY76, BM79, EMK82, RY77, vW80b]. sectors [ES77]. Secure [CM79b]. segmentation [vT72]. Selection [AN80, ST77, Ebe79, Erk80, Kit75, MZ82, Nos74, Rei78]. Self [Kru79]. Self-stabilization [Kru79]. semantic [CG76, RS77]. Semantics [Bro80b, GS76, EPS82, LNS82, MC80, Tal77]. semaphore [KM79, Tri76, Vv72]. semi [Rob79]. semilinear [Iba74]. Sensitivity [Tar82c, Tar86, KP72]. sentential [Roz72a]. Sequence [BWF^{+}79, GMS75, HG77]. Sequence-to-sequence [GMS75]. sequences [Ist78]. Sequential [Bud79, Fra79a, RM73, DZ71, GR76, GR77, Win77, Win78, ZD71]. series [Zav76]. Set [And78, Cza80, Jar73, SG76, Ban80, Byk78, DS79, Fou79, Fra79b, Gra72, Han76a, Han76b, Kay76, MZ79, Mat77b, Mis79, Ov80, SK77, SG79, YJ76]. Set-Theoretic [Cza80]. sets [Che79, ER73, Iba74, LR74, Rei78, Szw79]. several [FHMS79]. shallow [Pri79]. shared [CG78]. Sharing [CHCP74, Bv82, DS81, Dij72]. shop [Sło79]. Short [Fin72, Lyo79, Krá72a, Krá72b]. Shortest [Mor81, Rom80, NMM78, Tar82c, Tar86, Yuv76a]. Shortest-path [Mor81, Rom80]. Should [Kei75]. shuffle [Cah77]. shuffle-exchange [Cah77]. side [CG76]. signature [Lei80]. Significance [Dob79, HK79]. similarity [AT78a, AT79, Byk79]. similarly [GL80]. Simple [Bła77, DR76, Dwy73, GJPT78, BM82, Cam74, DDM^{+}75, Far82, Fre80, Han76c, MC80, MA79, Pău78, Rob75, Yuv78, Yuv79]. simpler [Mer78]. simplification [Der79]. simply [Mor79b]. Simulation [Laz75, UMS82, DZ71, Jav77, Jon77, ZD71]. Simulations [Gal76]. Single [AKS78]. sink [KST82]. sink-finding [KST82]. Size [AL75, GG82, Hoa74]. sizes [Wil80a]. Slices [DS78]. slicing [Roz76b]. Small [MOS76]. smoker [Red77]. Soft [Zim72]. Software [vT72, VL79, Han74]. Solution [BWF^{+}79, HM73, Loi79, MI73, dd76a, Akr79, AJV86, Bac82, CK77, Mau71, MN72, Mor79a, PS80, PS81, Rob75, Urb80]. Solutions [CHCP74, Hue78, Red77]. Solving [dd75, MS72]. Some [BN76, Bro79c, Cam78, Cul79, DS78, Gab76b, Gal76, Hos75, KA75, Kas73, Ko82, NI80, Nie75b, Tar82a, EO82, Gra79a, SS79, Ukk79, VL79]. Sort [Dob79, Dob80, HK79, Dob82]. Sorting [Dob78b, KS74a, KS74b, Kun77, MA80, Nog87, PE78, Bar73, Dew74, Dob78a, Erk80, JKS79, KK72, War73, Yuv73, Bur78]. Space [EL76, GI79, Mis79, Abe79, DLR79, KHTY78, Kaw80, Mee76, Meh76, Rob79, Yuv75, Zuc77, ten80, van77]. space-saving [Mee76]. Space-time [Mis79]. spacing [MS72]. Spanning [CC75, Cam78, Joh75, NM82, Roh78, Tar82c, Tar86, Yao75a, Yao75b]. specific [HV75]. specification [Cza79, PL79, Tal77]. specifications [BM82]. spider [Ebe79]. spite [Kru79]. split [KM79]. squares [Bac82, BMS80]. stability [Ber82, Day80b]. stabilization [Kru79]. stable [Bui77, Bui79, Day80a, Dew74, Kei75]. Stack [Dwy73, Pri79, Rob73, Sik72]. Stacking [Kar71]. Stacks [Lin73, May72]. star [Gab76a]. starvation [AJV86, DL78b, Mor79a]. starvation-free [AJV86, Mor79a]. Statements [DL73, KF71a, KF71b]. states [KD73]. Static [MG73]. Statistical [van79b, Yao79a]. statistics [KP72]. Stein [CM77]. steps [CLW80a, Sha79]. stiff [Bui77, Bui79, Day80a]. Stochastic [Bra72]. Storage [Bra72, CM79b, EL76, LP76, Rob80, Bro78, Bro79b, Cam74, Chr72, Dew74, Gel71, KHTY78, Kro73, LS82, Loz76, ten80]. store [Hoa74]. stores [Rem75]. straight [Dob82]. Strassen [Yuv79, Yuv78]. strategies [Bra78]. Strategy [Fin72, TW73]. Streams [CHCP74, Dij72]. String [Gal76, BD80]. Strong [GM82, Shi79c, FTM79, Tar82b]. strongly [BS79]. Structural [Chr76]. Structure [Gun82, BR75, CG78, KHTY78, Lao79]. Structured [GL75, KS74a, KS74b, Sol74, Zel74, GTY78, Kru79, Var72]. Structures [FW76b, FW79a, Lin73, All78, BB76a, CKH79]. study [BM77]. sub [Boa73]. sub-AFL [Boa73]. Subclass [FG78, Hug78]. subclasses [Sli82]. subdivision [LY79]. Subgraph [BB76a]. subgraphs [Gab76a, OO72]. subkeys [BD80]. subrecursive [Coy76, Tsi71]. Subsequence [Hir78]. Subsequences [BWF^{+}79]. subspaces [Kit77]. SUBSTAD [de 78]. Substantiable [PD79]. Substitution [de 78, Boa73]. substrings [Boo80]. subtext [BD80]. subwords [ER73]. sum [Bac82]. summations [Gon78, Gon79]. support [DDM^{+}75]. supporting [Ghe75]. swaps [Ita79b]. switching [PS80, PS81]. Symbolic [GL75, How78, CPR71]. symmetric [BES82]. symmetrizability [CM80]. symmetrizing [Col79, CM80]. Synchronization [DS78, DM79, Lam76, LT75]. Syntax [Coh73, BCDL72, EPS82]. Syntax-directed [Coh73]. syntax-free [EPS82]. Synthesis [KP72, Gin82]. System [Laz75, Loi79, RBE^{+}72, Sam76, BM77, BKH78, DD76b, FG79, Kar73, Kaw80, Lei80, LC76, Mad77, SY72, Tza72a, Var72, WB79, Zel74]. Systems [Bra72, Llo80b, MI73, Szw79, AR79, AR80, Ber82, CRM75a, CRM75b, Cul78, Cza78a, Gra80, Hug78, Ist78, Kob77, Kru79, Llo80a, Mer71, Pri76, Row75, Row77, Roz72a, Roz72c, Roz72b, Roz76a, Tsi72, Win77, Win78, Wür77, vT72, vW80a]. Szilard [Pău79].

T [Day80a]. T0L

[Roz72c, ER73, Roz72b, Von75].

T0L-systems [Roz72c, Roz72b]. Table [FW76a, FW77, San76]. Tabular [Gra79b]. Tag [Ist78, Lin73, Han76c]. Takeuchi [Moo79]. tale [LNS82]. Tape [SG77, FS73a]. Tarjan [Ban80]. Tasks [LF76, Bła79, Leu79, LM80a]. technique [Bar73, HN79, Jak82, Mee76]. Term [Fin72]. Terminals [Von75]. Termination [DS80, LLW79, Moo79]. tessellation [NI80]. test [Cel79, Che79, Len79]. Testing [Gav77, How78, Osb79, APT79, APT82, Bha80, DL78a, Gav75, LD80]. th [Urb80]. Theil [Yuv77]. their [PD79, ten80]. theorem [CRM75a, CRM75b, ER73, Kun79, YWL82, YWL85]. theorems [ES76, Kob77, LNS82, Paj80]. Theoretic [Cza80, Hir78]. theories [Kob74]. Theory [MG73, GTY78, MS72]. There [AKS78]. thirds [Gel71]. Three [BMT79, Ruo75, Bar76, IT80, YWL82, YWL85]. three-restricted [YWL82, YWL85]. three-way [IT80]. Thue [Ist78]. Time [Cof73, DR76, Gun82, PV80, UMS82, Abe79, APT79, APT82, BS78, FL82, FS73a, Fra79b, GT78, Gra77b, Gra77c, GL80, Han74, HJ71, Jav77, Jon77, Kaw80, KP72, LM80a, Mis79, PM75, Sli82, van77]. time-space [Abe79]. times [Bar73, Bła79]. tool [DY75, Sli82]. top [CKH79]. Topological [KS74a, KS74b, Yuv73]. total [Sou79]. Towers [BL80]. trade [Mis79]. tradeoffs [Abe79]. transducers [Eng78]. transfer [BM82]. Transferences [HM73]. transformation [Hen73]. Transformational [Bro80b]. transformations [Mat77b]. transformer [Fra77, MC80]. Transforms [Fey76, Ita79b]. Transition [Tsu78, Nie75a]. Transitive [Cof73, Mun71]. translations [Cul79]. translator [SY72]. transportable [WB79]. Traversal [Ada73]. Traversing [Dwy73, Mor79b, Rob73, Sik72]. Tree [Ada73, Cam78, Dwy73, NM82, Roh78, BN78a, BN78b, Bre80, CGH75, CRT79, Eng78, KA75, Kru79, Kun76, Kun77, Kun79, Rei79a, Sel77, Var72]. tree-structured [Kru79]. tree-to-tree [Sel77]. Trees [Bof73, CC75, Joh75, MOS76, Min73, NW78, NPWY72, Sam76, SF79, Wes76, Bay72, BS75, Bro78, Bro79b, Bro79c, Cla75, DKZ77, Eng72, FZ82, HR76, Kor82, MT75, MH79, Mor79b, Rob73, Ros79, Rot75, Sik72, Tar82c, Tar86, Weg75, Yao75a, Yao75b, Zak82]. Triangulating [GJPT78]. triangulation [MZ79]. tribute [Ano78]. Tries [Ore82]. truth [APT79, APT82]. Tsao [Bin82]. Tuning [LR80]. Turing [Dei74, SK77, SS77]. Two [Akl79, Bir77, Gal76, Har79a, Har79b, ITN78, KM79, Kob77, UMS82, And79, And80, AR79, Bac82, Byk78, CT73, DS79, ER79a, ER80, FS73a, Fou79, Fre80, Gel71, IT80, Ita79b, Jon77, NI80, Ov80, Paj80, Ple79, Row75, Row77]. Two-Dimensional [ITN78, Fre80, IT80, NI80]. two-processor [AR79, Row75, Row77]. two-thirds [Gel71]. Two-Way [UMS82, Jon77]. Type [Gun82, SG77, Ber78, FG79, Kar73, PL79, Rot75]. typed [BS79]. Types [Han75, BS79, Lau80].

UET [Llo80a, vW80a, Leu79, Llo80b]. Ullman [Yao79a]. unbounded [BY76]. undecidability [Roz72a]. undecidable [Roz72c, Roz72b]. Understanding [PR77]. undirected [Ebe80]. unicyclic [MH79]. Unification [de 78]. uniformly [LR74]. unilaterally [Arj76, Che78]. unintentional [Bak72]. Union [DR76, Ban80, Lao79]. Union-Find [DR76, Lao79]. unique [MN72]. Unit [LF76, Coh73]. Unit-Length [LF76]. United [BS79]. Universal [Han75, Her73, Zuc77]. universally [SK77]. unpredictability [HB77]. unravelling [DY75]. unsolvability [Gra79a]. unstructured [Oul79]. Update [Joh75]. upper [Kun73, Nie75b]. use [Cam74, OO72, PL79]. Used [Arn74, Ore82, Var72]. Users [Sam76]. Using [CM79b, GW78, HG77, Mak79, MOS76, Akr79, Han74, Hor72a, Hor72b, KM79, Sch78, Tar82b, Tes79, Yuv76a].

V [SG77]. V-Tape [SG77]. valuation [Kun79]. value [Kun79, Wha72]. Values [Gra77a, Gra79b, Mee76]. Variables [MAP76, HJ71]. variant [Dei74, Han76c]. Variation [Dob80]. Variational [dd76a]. Vector [Kam75, MRK76, Avi80, CRM75a, CRM75b, DV76, Gra80, Hug78, Jak78, Jak82, Jan79, van80]. vector-addition [CRM75a, CRM75b]. vector-products [DV76]. Vector/Parallel [MRK76]. vectors [Dev80, Teu78]. verification [PL79]. Verifier [GL75]. verifying [WS82]. versus [Roz76a]. Vertex [Sri74, Sri75, SG79]. vertices [GT78]. Very [MOS76]. via [Dev80, Han80, LR80]. Virtual [SG77, Sta71]. VLSI [Joh80, PV80]. Voronoi [Bro79a].

Waite [Der80]. Warren [Dob78a]. Way [UMS82, IT80, Jon77]. weak [Sud77, Sud78, Tar74c]. Weighted [NW78, NPWY72]. Weights [MTT73, Rei78]. whether [Tar72]. Which [Bof73, FW76a, FW77, Gal76, Tsu78]. Within [BWF^{+}79, Var72]. Without [Lin73, Bro72, DL73, Dwy73, Kar71, Red77, Rob73, Ruo75, Sik72]. word [Krá72a, Krá72b]. word-length [Krá72a, Krá72b]. words [ER79a]. working [Chr72]. Workspace [Fis74]. worst [Röh82]. writer [BN76]. writers [BN78a, BN78b]. Writing [Bac82, FW76b, FW79a].

Xcribl [RBE^{+}72].

References

Abelson:1979:NTS

[Abe79] Harold Abelson. A note on timespace tradeoffs for computing continuous functions. Information Processing Letters, 8(4):215–217, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Adams:1973:ARB

[Ada73] Edward N. Adams, III. Another representation of binary tree traversal. Information Processing Letters, 2(2):52–54, June ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderssen:1975:RND

[AG75] R. S. Anderssen and A. J. Guttmann. A rationale for the numerical differentiation of experimental data. Information Processing Letters, 4(2):48–50, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Annot:1986:CMS

[AJV86] J. K. Annot, M. D. Janssens, and A. J. Van De Goor. Comments on Morris’s starvation-free solution to the mutual exclusion problem. Information Processing Letters, 23 (2):91–97, August 20, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mor79a].

Akl:1978:CGM

[Akl78] S. G. Akl. Comments on: G. Manacher, “An application of pattern matching to a problem in geometrical complexity”. Information Processing Letters, 7(2):86–??, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akl:1979:TRC

[Akl79] S. G. Akl. Two remarks on a convex hull algorithm. Information Processing Letters, 8(2):108–109, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akritas:1979:SPE

[Akr79] A. G. Akritas. On the solution of polynomial equations using continued fractions. Information Processing Letters, 9(4):182–184, November 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ajtai:1978:TNF

[AKS78] Miklós Ajtai, János Komlós, and Endre Szemerédi. There is no fast single hashing algorithm. Information Processing Letters, 7(6):270– 273, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Asai:1975:DQB

[AL75] H. Asai and S. C. Lee. Design of queueing buffer register size. Information Processing Letters, 3(5): 147–152, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1978:PSN

[All78] L. Allison. Phrase structures, nondeterminism and backtracking. Information Processing Letters, 7(3): 139–143, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albert:1977:CCF

[AM77] J. Albert and H. A. Maurer. The class of context-free languages is not an E0L family. Information Processing Letters, 6(6):190–195, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1980:SDP

[AN80] D. C. S. Allison and M. T. Noga. Selection by distributive partitioning. Information Processing Letters, 11 (1):7–8, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderssen:1972:RPP

[And72] R. S. Anderssen. A refinement procedure for pure random search. Information Processing Letters, 1(5): 197–200, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderson:1978:REA

[And78] Kenneth R. Anderson. A reevaluation of an efficient algorithm for determining the convex hull of a finite planar set. Information Processing Letters, 7(1):53–55, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andrew:1979:AEA

[And79] A. M. Andrew. Another efficient algorithm for convex hulls in two dimensions. Information Processing Letters, 9(5):216–219, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [And80].

Andrew:1980:CAE

[And80] A. M. Andrew. Corrigendum: “Another efficient algorithm for convex hulls in two dimensions” (Inform. Process. Lett. 9 (1979), no. 5, 216–219). Information Processing Letters, 10(3):168, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [And79].

Angluin:1979:NCM

[Ang79] Dana Angluin. A note on a construction of Margulis. Information Processing Letters, 8(1):17–19, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:1978:TR

[Ano78] Anonymous. A tribute to referees. Information Processing Letters, 7(1):64, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aspvall:1979:LTA

[APT79] Bengt Aspvall, Michael F. Plass, and Robert Endre Tarjan. A lineartime algorithm for testing the truth of certain quantified Boolean formulas. Information Processing Letters, 8(3):121–123, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [APT82].

Aspvall:1982:ELT

[APT82] Bengt Aspvall, Michael F. Plass, and Robert Endre Tarjan. Erratum: “A linear-time algorithm for testing the truth of certain quantified Boolean formulas” [Inform. Process. Lett. 8 (1979), no. 3, 121–123; MR 80b:68050]. Information Processing Letters, 14(4):195, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [APT79].

Arora:1979:MAT

[AR79] R. K. Arora and S. P. Rana. On module assignment in twoprocessor distributed systems. Information Processing Letters, 9(3): 113–117, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1980:HAP

[AR80] R. K. Arora and S. P. Rana. Heuristic algorithms for process assignment in distributed computing systems. Information Processing Letters, 11(4–5):199–203, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arjomandi:1976:FAU

[Arj76] E. Arjomandi. On finding all unilaterally connected components of a digraph. Information Processing Letters, 5(1):8–10, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arnborg:1974:ACM

[Arn74] Stefan Arnborg. Abstract computation model used for a production compiler. Information Processing Letters, 3(2):46–50, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1978:EOP

[AS78a] Radha Krishan Arora and R. K.

Subramanian. Exploiting the optimal paging algorithms. Information Processing Letters, 7(5):233– 236, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1978:ODP

[AS78b] Radha Krishan Arora and R. K.

Subramanian. An optimal demand prepaging algorithm. Information Processing Letters, 7(3):132– 136, April 28, 1978. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akl:1978:IAC

[AT78a] S. G. Akl and G. T. Toussaint.

An improved algorithm to check for polygon similarity. Information Processing Letters, 7(3):127–128, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See addendum [AT79].

Akl:1978:FCH

[AT78b] Selim G. Akl and Godfried T. Toussaint. A fast convex hull algorithm. Information Processing Letters, 7(5):219–222, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akl:1979:AIA

[AT79] S. G. Akl and G. T. Toussaint. Addendum to “An improved algorithm to check for polygon similarity”. Information Processing Letters, 8(3): 157–158, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [AT78a].

Avis:1980:CLB

[Avi80] David Avis. Comments on a lower bound for convex hull determination: “On the Ω(n log n) lower bound for convex hull and maximal vector determination” by van Emde Boas [Inform. Process. Lett. 10(3), 18 April 1980, pp. 132–136]. Information Processing Letters, 11(3): 126, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [van80].

Backhouse:1982:WNS

[Bac82] Ronald Backhouse. Writing a number as the sum of two squares: a new solution. Information Processing Letters, 14(1):15–17, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baker:1972:UOA

[Bak72] J. L. Baker. An unintentional omission from ALGOL 68. Information Processing Letters, 1(6):244– 245, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Banatre:1977:POC

[Ban77] J. P. Banatre. Producing optimised code for coercions. Information Processing Letters, 6(2):56–59, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Banachowski:1980:CTR

[Ban80] Lech Banachowski. A complement to Tarjan’s result about the lower bound on the complexity of the set union problem. Information Processing Letters, 11(2):59–65, October ??, 1980. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Barnett:1973:TRC

[Bar73] J. K. R. Barnett. A technique for reducing comparison times in certain applications of the merging method of sorting. Information Processing Letters, 2(5):127–128, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barak:1976:PED

[Bar76] A. B. Barak. On the parallel evaluation of division-free arithmetic expressions with fan-in of three. Information Processing Letters, 5(1):18– 19, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bayer:1972:OBT

[Bay72] R. Bayer. Oriented balanced trees and equivalence relations. Information Processing Letters, 1(6):226– 228, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barrow:1976:SIM

[BB76a] H. G. Barrow and R. M. Burstall.

Subgraph isomorphism, matching relational structures and maximal cliques. Information Processing Letters, 4(4):83–84, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1976:HPM

[BB76b] Jon Louis Bentley and Walter A.

Burkhard. Heuristics for partialmatch retrieval data base design. Information Processing Letters, 4(5):

132–135, February ??, 1976. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonuccelli:1979:MND

[BB79] M. A. Bonuccelli and D. P. Bovet. Minimum node disjoint path covering for circular-arc graphs. Information Processing Letters, 8(4): 159–161, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bini:1979:LBC

[BC79] Dario Bini and Milvio Capovani. Lower bounds of the complexity of linear algebras. Information Processing Letters, 9(1):46–47, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Branquart:1972:CFS

[BCDL72] P. Branquart, J. P. Cardinael, J. P. Delescaille, and J. Lewi. A contextfree syntax of ALGOL 68. Information Processing Letters, 1(4): 141–148, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bini:1979:CAM

[BCRL79] Dario Bini, Milvio Capovani, Francesco Romani, and Grazia Lotti. O(n^{2.7799}) complexity for n × n approximate matrix multiplication. Information Processing Letters, 8(5):234–235, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bailey:1980:FSS

[BD80] T. A. Bailey and R. G. Dromey. Fast string searching by finding subkeys in subtext. Information Processing Letters, 11(3):130–133, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1979:DSP

[Ben79] Jon Louis Bentley. Decomposable searching problems. Information Processing Letters, 8(5):244– 251, June 11, 1979. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bergstra:1978:WAD

[Ber78] Jan Bergstra. What is an abstract data type? Information Processing Letters, 7(1):42–43, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beretta:1982:MCE

[Ber82] G. Beretta. Monte Carlo estimation of numerical stability in fast algorithms for systems of bilinear forms. Information Processing Letters, 14 (4):162–167, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Best:1979:NPC

[Bes79] Eike Best. A note on the proof of a concurrent program. Information Processing Letters, 9(3):103– 104, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barlow:1982:PMD

[BES82] R. H. Barlow, D. J. Evans, and J. Shanehchi. Parallel multisection for the determination of the eigenvalues of symmetric quindiagonal matrices. Information Processing Letters, 14(3):117–118, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhat:1980:CTG

[Bha80] Kabekode V. S. Bhat. On the complexity of testing a graph for n-cube. Information Processing Letters, 11 (1):16–19, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bini:1982:RPN

[Bin82] Dario Bini. Reply to the paper: “The numerical instability of Bini’s algorithm” [Inform. Process. Lett. 12 (1981), no. 1, 17–19; MR 82i:65029] by N. K. Tsao. Information Processing Letters, 14(3):144– 145, May 16, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Tsa81].

Bird:1977:TDP

[Bir77] R. S. Bird. Two dimensional pattern matching. Information Processing Letters, 6(5):168–170, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandwajn:1978:ASR

[BKH78] A. Brandwajn, Ph. Kruchten, and J. A. Hernandez. ARCADE: a system for research and education in computer architecture. Information Processing Letters, 7(2):78– 85, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bongiovanni:1978:CRG

[BL78] G. Bongiovanni and F. Luccio. On Cahit’s result on graceful permutations. Information Processing Letters, 7(4):183–184, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buneman:1980:THP

[BL80] Peter Buneman and Leon Levy. The Towers of Hanoi problem. Information Processing Letters, 10(4–5): 243–244, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blazewicz:1977:SAM

[Bła77] J. Błazewicz. Simple algorithms for multiprocessor scheduling to meet deadlines. Information Processing Letters, 6(5):162–164, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blazewicz:1979:DST

[Bła79] Jacek Błażewicz. Deadline scheduling of tasks with ready times and resource constraints. Information Processing Letters, 8(2):60–63, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borodin:1971:EPM

[BM71] A. Borodin and I. Munro. Evaluating polynomials at many points. Information Processing Letters, 1 (2):66–68, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandwajn:1977:SPD

[BM77] A. Brandwajn and B. Mouneix. A study of a page-on-demand system. Information Processing Letters, 6 (4):125–132, August ??, 1977. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1979:NEN

[BM79] Jon Louis Bentley and Hermann A. Maurer. A note on Euclidean near neighbor searching in the plane. Information Processing Letters, 8(3): 133–136, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bergstra:1982:STL

[BM82] J. A. Bergstra and J.-J. Ch. Meyer. A simple transfer lemma for algebraic specifications. Information Processing Letters, 14(2):80–85, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bohm:1980:CBE

[BMS80] Corrado Böhm, Antonio Machì, and Giovanna Sontacchi. Complexity bounds for equivalence and isomorphism of Latin squares. Information Processing Letters, 10(4–5):231–233, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertoni:1979:TEA

[BMT79] Alberto Bertoni, Giancarlo Mauri, and Mauro Torelli. Three efficient algorithms for counting problems. Information Processing Letters, 8 (1):50–53, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boari:1976:SPD

[BN76] M. Boari and A. Natali. Some properties of deadlock detection and recovery in readers’ and writers’ problems. Information Processing Letters, 5(4):118–123 (or 118–122??), October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boari:1978:EMA

[BN78a] M. Boari and A. Natali. Erratum:

“Multiple access to a tree in the context of readers and writers problem”. Information Processing Letters, 7(4):207, June ??, 1978. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BN78b].

Boari:1978:MAT

[BN78b] M. Boari and A. Natali. Multiple access to a tree in the context of readers and writers problem. Information Processing Letters, 7(2):112– 121, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [BN78a].

Boasson:1973:ISC

[Boa73] L. Boasson. The inclusion of the substitution closure of linear and onecounter languages in the largest sub-AFL of the family of algebraic languages is proper. Information Processing Letters, 2(5):135–140, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bochmann:1976:CMD

[Boc76] G. V. Bochmann. Comments on monitor definition and implementation. Information Processing Letters, 5(4):116–117, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boffey:1973:AMR

[Bof73] T. B. Boffey. Applying the minimax rule over graphs which are not trees. Information Processing Letters, 2(3):79–81, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Booth:1980:LLC

[Boo80] Kellogg S. Booth. Lexicographically least circular substrings. Information Processing Letters, 10(4– 5):240–242, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Banerjee:1975:DLD

[BR75] J. Banerjee and V. Rajaraman. A dual link data structure for random file organization. Information Processing Letters, 4(3):64–69, December ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandstetter:1972:SRS

[Bra72] A. Brandstetter. Storage requirements in stochastic data acquisition systems. Information Processing Letters, 1(6):217–219, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bratko:1978:PCS

[Bra78] I. Bratko. Proving correctness of strategies in the AL1 assertional language. Information Processing Letters, 7(5):223–230, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brent:1980:ABT

[Bre80] R. P. Brent. On the area of binary tree layouts. Information Processing Letters, 11(1):46–48, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bron:1972:OMB

[Bro72] C. Bron. Outline of a machine without branch instructions. Information Processing Letters, 1(3):117– 119, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brown:1978:SSH

[Bro78] M. R. Brown. A storage scheme for height-balanced trees. Information Processing Letters, 7(5):231– 232, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See addendum [Bro79b].

Brown:1979:VDC

[Bro79a] Kevin Q. Brown. Voronoi diagrams from convex hulls. Information Processing Letters, 9(5):223–228, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brown:1979:ASS

[Bro79b] M. R. Brown. Addendum to “A storage scheme for height-balanced trees”. Information Processing Letters, 8(3):154–156, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bro78].

Brown:1979:SOR

[Bro79c] M. R. Brown. Some observations on random 2-3 trees. Information Processing Letters, 9(2):57–59, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brown:1980:IBL

[Bro80a] Donna J. Brown. An improved BL lower bound. Information Processing Letters, 11(1):37–39, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broy:1980:TSC

[Bro80b] M. Broy. Transformational semantics for concurrent programs. Information Processing Letters, 11(2): 87–91, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1975:ARS

[BS75] J. L. Bentley and D. F. Stanat. Analysis of range searches in quad trees. Information Processing Letters, 3(6):170–173, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1978:DCL

[BS78] J. L. Bentley and M. I. Shamos. Divide and conquer for linear expected time. Information Processing Letters, 7(2):87–91, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berry:1979:UDR

[BS79] D. M. Berry and R. L. Schwartz.

United and discriminated record types in strongly typed languages. Information Processing Letters, 9 (1):13–18, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1977:CFF

[BSW77] Jon L. Bentley, Donald F. Stanat, and E. Hollins Williams, Jr. The complexity of finding fixed-radius near neighbors. Information Processing Letters, 6(6):209–212, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buda:1979:GSM

[Bud79] A. Buda. Generalized ^{1,5} sequential machine maps. Information Processing Letters, 8(1):38–40, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bui:1977:SMS

[Bui77] T. D. Bui. On an L-stable method for stiff differential equations. Information Processing Letters, 6(5): 158–161, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Bui79] and comments [Day80a].

Bui:1979:ESM

[Bui79] T. D. Bui. Erratum: “On an Lstable method for stiff differential equations”. Information Processing Letters, 8(4):218, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bui77, Day80a].

Burton:1978:CSD

[Bur78] W. Burton. Comments on: Sorting by distributive partitioning. Information Processing Letters, 7(4): 205–206, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Dob78b, Dob78a, HK79].

Brouwer:1982:NMK

[Bv82] Andries E. Brouwer and Peter van Emde Boas. A note on: “Master keys for group sharing” [Inform. Process. Lett. 12 (1981), no. 1, 23– 25; MR 82d:94046] by D. E. Denning and F. B. Schneider. Information Processing Letters, 14(1):12–14, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [DS81].

Broy:1979:MSP

[BWF^{+}79] M. Broy, M. Wirsing, J.-P. Finance, A. Quéré, and J.-L. Remy. Methodical solution of the problem of ascending subsequences of maximum length within a given sequence. Information Processing Letters, 8 (5):224–229, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bentley:1976:AOA

[BY76] J. L. Bentley and A. Chi-Chih Yao.

An almost optimal algorithm for unbounded searching. Information Processing Letters, 5(3):82–87, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bykat:1978:CHF

[Byk78] A. Bykat. Convex hull of a finite set of points in two dimensions. Information Processing Letters, 7(6): 296–298, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [DS79, Fou79, Ov80].

Bykat:1979:PS

[Byk79] A. Bykat. On polygon similarity. Information Processing Letters, 9(1): 23–25, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cahit:1977:RGP

[Cah77] I. Cahit. Realization of graceful permutation by a shuffle-exchange network. Information Processing Letters, 6(5):171–173, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Campbell:1974:OUS

[Cam74] J. A. Campbell. Optimal use of storage in a simple model of garbage collection. Information Processing Letters, 3(2):37–38, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Camerini:1978:MMS

[Cam78] P. M. Camerini. The min-max spanning tree problem and some extensions. Information Processing Letters, 7(1):10–14, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cahit:1975:GNS

[CC75] I. Cahit and R. Cahit. On the graceful numbering of spanning trees. Information Processing Letters, 3(4): 115–118, March ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caviness:1978:NCA

[CE78] B. F. Caviness and H. I. Epstein. A note on the complexity of algebraic differentiation. Information Processing Letters, 7(2):122–124, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cellary:1979:NST

[Cel79] Wojciech Cellary. A new safety test for deadlock avoidance. Information Processing Letters, 8(3): 115–120, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carroll:1972:ACA

[CFG72] J. M. Carroll, W. Fraser, and G. Gill. Automatic content analysis in an on-line environment. Information Processing Letters, 1(4): 134–140, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cunningham:1976:NSD

[CG76] R. J. Cunningham and M. E. J. Gilford. A note on the semantic definition of side effects. Information Processing Letters, 4(5):118–120, February ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clark:1978:NSL

[CG78] D. W. Clark and C. C. Green.

A note on shared list structure in LISP. Information Processing Letters, 7(6):312–314, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cockayne:1975:LAD

[CGH75] E. Cockayne, S. Goodman, and S. Hedetniemi. A linear algorithm for the domination number of a tree. Information Processing Letters, 4(2):41–44, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cooprider:1974:ISS

[CHCP74] Lee W. Cooprider, F. Heymans, P. J. Courtois, and David L. Parnas. Information streams sharing a finite buffer: Other solutions. Information Processing Letters, 3(1):16–21, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheston:1978:CUC

[Che78] G. A. Cheston. A correction to a unilaterally connected components algorithm. Information Processing Letters, 7(2):125–??, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cherniavsky:1979:FTD

[Che79] J. C. Cherniavsky. On finding test data sets for loop free programs. Information Processing Letters, 8(2): 106–107, February 15, 1979. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chor:1982:AFF

[Cho82] Ben-Zion Chor. Arithmetic of finite fields. Information Processing Letters, 14(1):4–6, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chroust:1972:EEM

[Chr72] G. Chroust. Expression evaluation with minimum average working storage. Information Processing Letters, 1(3):111–114, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Christensen:1975:CME

[Chr75] R. Christensen. Crossvalidation: Minimizing the entropy of the future. Information Processing Letters, 4(3):73–76, December ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobot:1976:LLC

[Chr76] Stanislaw Chrobot. Layer — a language construction for concurrent structural program design. Information Processing Letters, 4(5):113– 117, February ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1977:ASC

[CK77] J. Cohen and J. Katcoff. Automatic solution of a certain class of combinatorial problems. Information Processing Letters, 6(4):101–104, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cherniavsky:1979:NCT

[CKH79] John C. Cherniavsky, John Keohane, and Peter B. Henderson. A note concerning top down program development and restricted exit control structures. Information Processing Letters, 9(1):8–12, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calmet:1980:IRP

[CL80] J. Calmet and R. Loos. An improvement of Rabin’s probabilistic algorithm for generating irreducible polynomials over GF(p). Information Processing Letters, 11(2):94–95, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clark:1975:FAC

[Cla75] D. W. Clark. A fast algorithm for copying binary trees. Information Processing Letters, 4(3):62–63, December ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1980:MNS

[CLW80a] K. M. Chung, F. Luccio, and C. K.

Wong. Minimum number of steps for permutation in a bubble memory. Information Processing Letters, 11(2):81–83, October ??, 1980.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1980:NPA

[CLW80b] K. M. Chung, F. Luccio, and C. K. Wong. A new permutation algorithm for bubble memories. Information Processing Letters, 10(4–5): 226–230, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Camerini:1975:BMI

[CM75] P. M. Camerini and F. Maffioli. Bounds for 3-matroid intersection problems. Information Processing Letters, 3(3):81–83, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Collins:1977:APS

[CM77] George E. Collins and David R. Musser. Analysis of the Pope– Stein division algorithm. Information Processing Letters, 6(5):151– 155, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandy:1979:DAP

[CM79a] K. M. Chandy and J. Misra. Deadlock absence proofs for networks of communicating processes. Information Processing Letters, 9(4):185– 189, November 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1979:SIS

[CM79b] K. Culik, II and H. A. Maurer. Secure information storage and retrieval using new results in cryptography. Information Processing Letters, 8(4):181–186, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colbourn:1980:CCP

[CM80] Charles J. Colbourn and Brendan D.

McKay. A correction to Colbourn’s paper on the complexity of matrix symmetrizability: “The complexity of symmetrizing matrices” [Inform. Process. Lett. 9 (1979), no. 3, 108– 109; MR 81a:68045]. Information Processing Letters, 11(2):96–97, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Col79].

Cremers:1973:NLR

[CMM73] A. B. Cremers, H. A. Maurer, and O. Mayer. A note on leftmost restricted random context grammars. Information Processing Letters, 2 (2):31–33, June ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Coffy:1973:CTC

[Cof73] Joel Coffy. On computing the time complexity of transitive closure algorithms. Information Processing Letters, 2(2):39–42, June ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1973:SDU

[Coh73] J. Cohen. Syntax-directed unit conversion. Information Processing Letters, 2(4):100–102, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1975:IND

[Coh75] Jacques Cohen. Interpretation of non-deterministic algorithms in higher-level languages. Information Processing Letters, 3(4):104– 109, March ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colbourn:1979:CSM

[Col79] Charles J. Colbourn. The complexity of symmetrizing matrices. Information Processing Letters, 9(3): 108–109, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also correction [CM80].

Corneil:1971:ADB

[Cor71] D. G. Corneil. An n^{2} algorithm for determining the bridges of a graph. Information Processing Letters, 1 (2):51–55, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Coy:1976:LMP

[Coy76] Wolfgang Coy. The logical meaning of programs of a subrecursive language. Information Processing Letters, 4(5):121–126, February ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caviness:1971:ELC

[CPR71] B. F. Caviness, P. L. Pollack, and C. M. Rubald. An existence lemma for canonical forms in symbolic mathematics. Information Processing Letters, 1(2):45–46, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crespi-Reghizzi:1975:DTC

[CRM75a] S. Crespi-Reghizzi and D. Mandrioli. A decidability theorem for a class of vector-addition systems. Information Processing Letters, 3(3): 78–80, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [CRM75b].

Crespi-Reghizzi:1975:EDT

[CRM75b] S. Crespi-Reghizzi and D. Mandrioli. Erratum: “A decidability theorem for a class of vector-addition systems”. Information Processing Letters, 3(5):164, May ??, 1975. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CRM75a].

Crespi-Reghizzi:1978:CGG

[CRM78] Stefano Crespi-Reghizzi and Dino Mandrioli. A class of grammar generating non-counting languages. Information Processing Letters, 7(1): 24–26, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cockayne:1979:AME

[CRT79] E. J. Cockayne, F. Ruskey, and A. G. Thomason. An algorithm for the most economic link addition in a tree communications network. Information Processing Letters, 9(4):171– 175, November 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:1977:FEE

[CS77] F. Chin and K. Steiglitz. A fast error evaluation algorithm for polynomial approximation. Information Processing Letters, 6(1):18–21, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Coffman:1973:NRP

[CT73] E. G. Coffman, Jr. and C. J. M. Turnbull. A note on the relative performance of two disk scanning policies. Information Processing Letters, 2(1):15–17, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1978:DLC

[Cul78] Karel Culik, II. The decidability of ν-local catenativity and of other properties of D0L systems. Information Processing Letters, 7(1):33–35, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1979:SDR

[Cul79] K. Culik, II. Some decidability results about regular and pushdown translations. Information Processing Letters, 8(1):5–8, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1978:IAP

[Cza78a] L. Czaja. Implementation approach to parallel systems. Information Processing Letters, 7(5):244– 249, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1978:PIP

[Cza78b] Ludwik Czaja. Parallel implementation of path expressions. Information Processing Letters, 7(6):291– 295, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1979:SPP

[Cza79] Ludwik Czaja. A specification of parallel problems. Information Processing Letters, 8(4):162–167, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1980:DFP

[Cza80] Ludwik Czaja. Deadlock and fairness in parallel schemas: a settheoretic characterization and decision problems. Information Processing Letters, 10(4–5):234–239, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Day:1980:CSM

[Day80a] J. D. Day. Comments on: “On an L-stable method for stiff differential equations” [Inform. Process. Lett. 6 (1977), no. 5, 158–161; MR 56 #10015] by T. D. Bui. Information Processing Letters, 11(1):31–32, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bui77, Bui79].

Day:1980:ISR

[Day80b] J. D. Day. On the internal Sstability of Rosenbrock methods. Information Processing Letters, 11(1): 27–30, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deWolf:1975:NMS

[dd75] S. de Wolf and G. de Mey. Numerical methods for solving integral equations of potential problems. Information Processing Letters, 3(4): 121–124, March ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deWolf:1976:NSI

[dd76a] S. de Wolf and G. de Mey. Numerical solution of integral equations for potential problems by a variational principle. Information Processing Letters, 4(5):136–139, February ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Delcour:1976:EAC

[DD76b] K. Delcour and A. J. W. Duijvestein. Enclosures: an access control mechanism with applications in parallel programming and other areas of system programming. Information Processing Letters, 5(5):125– 135, November ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DalBono:1975:SFS

[DDM^{+}75] Ivan Dal Bono, Mauro Diligenti, Concetta Mosca, Antonio Ricci, and Antonio Villani. A simple FORTRAN support for computerassisted instruction. Information Processing Letters, 3(3):88–90, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deGroote:1975:CQM

[de 75] H. F. de Groote. On the complexity of quaternion multiplication. Information Processing Letters, 3(6): 177–179, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deChampeaux:1978:SFS

[de 78] Dennis de Champeaux. SUBSTAD:

For fast substitution in LISP, with an application on unification. Information Processing Letters, 7(1): 58–62, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deimel:1974:RCP

[Dei74] L. E. Deimel, Jr. Remark on the computational power of a Turing machine variant. Information Processing Letters, 3(2):43–45, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demetrovics:1978:NCK

[Dem78] János Demetrovics. On the number of candidate keys. Information Processing Letters, 7(6):266–269, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demolombe:1982:GDR

[Dem82] R. Demolombe. Generalized division for relational algebraic language. Information Processing Letters, 14(4): 174–178, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dershowitz:1979:NSO

[Der79] Nachum Dershowitz. A note on simplification orderings. Information Processing Letters, 9(5):212– 215, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dershowitz:1980:SMA

[Der80] N. Dershowitz. The Schorr–Waite marking algorithm revisited. Information Processing Letters, 11(3): 141–143, November 18, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devroye:1980:NFC

[Dev80] Luc Devroye. A note on finding convex hulls via maximal vectors. Information Processing Letters, 11(1): 53–56, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dewar:1974:SMS

[Dew74] Robert B. K. Dewar. A stable minimum storage sorting algorithm. Information Processing Letters, 2(6): 162–164, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1972:ISS

[Dij72] E. W. Dijkstra. Information streams sharing a finite buffer. Information Processing Letters, 1(5):179– 180, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deo:1977:GLG

[DKP77] Narsingh Deo, M. S. Krishnamoorthy, and Ajit B. Pai. Generalizations of line graphs and applications. Information Processing Letters, 6(1): 14–17, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dinic:1977:NTS

[DKZ77] E. A. Dinic, A. K. Kelmans, and M. A. Zaitsev. Nonisomorphic trees with the same T -polynomial. Information Processing Letters, 6(3): 73–76, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devillers:1973:RPN

[DL73] R. Devillers and G. Louchard. Realization of Petri nets without conditional statements. Information Processing Letters, 2(4):105–107, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeMillo:1978:PRA

[DL78a] R. A. DeMillo and R. J. Lipton.

A probabilistic remark on algebraic program testing. Information Processing Letters, 7(4):193–195, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devillers:1978:GMA

[DL78b] R. E. Devillers and P. E. Lauer. A general mechanism for avoiding starvation with distributed control. Information Processing Letters, 7(3): 156–158, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobkin:1979:LPL

[DLR79] David Dobkin, Richard J. Lipton, and Steven P. Reiss. Linear programming is log-space hard for P.

Information Processing Letters, 8 (2):96–97, February 15, 1979. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeMillo:1979:ICS

[DM79] R. A. De Millo and R. E. Miller. Implicit computation of synchronization primitives. Information Processing Letters, 9(1):35–38, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobosiewicz:1978:ARW

[Dob78a] W. Dobosiewicz. Author’s reply to Warren Burton’s comments on distributive partitioning sorting. Information Processing Letters, 7(4):206, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bur78, HK79].

Dobosiewicz:1978:SDP

[Dob78b] Włodzimierz Dobosiewicz. Sorting by distributive partitioning. Information Processing Letters, 7(1):1–6, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL https://www.

sciencedirect.com/science/article/
abs/pii/0020019078900285. See

debate [Bur78, HK79, Dob79, JKS79, Nog87].

Dobosiewicz:1979:PSD

[Dob79] Włodzimierz Dobosiewicz. The practical significance of D.P. sort revisited. Information Processing Letters, 8(4):170–172, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). This paper rebuts comments [HK79] on the original paper [Dob78b].

Dobosiewicz:1980:EVB

[Dob80] Włodzimierz Dobosiewicz. An efficient variation of bubble sort. Information Processing Letters, 11(1): 5–6, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doberkat:1982:AEH

[Dob82] Ernst-E. E. Doberkat. Asymptotic estimates for the higher moments of the expected behavior of straight insertion sort. Information Processing Letters, 14(4):179–182, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duske:1976:PRI

[DPS76] J. Duske, R. Parchmann, and H. Schumacher. A pattern representation of indexed languages. Information Processing Letters, 5(5):152– 154, November ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doyle:1976:LET

[DR76] Jon Doyle and Ronald L. Rivest.

Linear expected time of a simple union-find algorithm. Information Processing Letters, 5(5):146– 148, November ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dolev:1978:CRS

[DS78] Danny Dolev and Eliahu Shamir.

Commutation relations of slices characterize some synchronization primitives. Information Processing Letters, 7(1):7–9, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devai:1979:CCH

[DS79] F. Dévai and T. Szendrényi. Comments on convex hull of a finite set of points in two dimensions. Information Processing Letters, 9(3): 141–142, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Byk78, Fou79, Ov80].

Dijkstra:1980:TDD

[DS80] Edsger W. Dijkstra and C. S. Scholten. Termination detection for diffusing computations. Information Processing Letters, 11(1):1–4, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Denning:1981:MKG

[DS81] Dorothy E. Denning and Fred B. Schneider. Master keys for group sharing. Information Processing Letters, 12(1):23–25, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also note [Bv82].

Dasgupta:1975:MCM

[DT75] Subrata Dasgupta and John Tartar. On the minimization of control memories. Information Processing Letters, 3(3):71–74, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobkin:1976:CV

[DV76] D. Dobkin and J. Van Leeuwen. The complexity of vector-products. Information Processing Letters, 4(6): 149–154, March ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dwyer:1973:SAT

[Dwy73] Barry Dwyer. Simple algorithms for traversing a tree without an auxiliary stack. Information Processing Letters, 2(5):143–145, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duncan:1975:LUP

[DY75] A. G. Duncan and L. Yelowitz. Loop unravelling: a practical tool in proving program correctness. Information Processing Letters, 4(3):70–72, December ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dyer:1980:FPA

[Dye80] Charles R. Dyer. A fast parallel algorithm for the closest pair problem. Information Processing Letters, 11 (1):49–52, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duncan:1971:PSS

[DZ71] F. G. Duncan and D. Zissos. Programmed simulation of sequential circuits. Information Processing Letters, 1(1):4–6, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eberlein:1979:NMS

[Ebe79] P. J. Eberlein. A note on median selection and spider production. Information Processing Letters, 9(1):19– 22, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ebert:1980:NOE

[Ebe80] Jürgen Ebert. A note on odd and even factors of undirected graphs. Information Processing Letters, 11 (2):70–72, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrich:1976:SSR

[EL76] H. D. Ehrich and W. Lipski, Jr. On the storage space requirement of consecutive retrieval with redundancy. Information Processing Letters, 4(4):101–104, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1982:PIS

[EMK82] H. Edelsbrunner, H. A. Maurer, and D. G. Kirkpatrick. Polygonal intersection searching. Information Processing Letters, 14(2):74–79, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:1972:NIT

[Eng72] J. Engelfriet. A note on infinite trees. Information Processing Letters, 1(6):229–232, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:1978:TTP

[Eng78] Joost Engelfriet. On tree transducers for partial functions. Information Processing Letters, 7(4): 170–172, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1982:ESR

[EO82] Herbert Edelsbrunner and Mark H.

Overmars. On the equivalence of some rectangle problems. Information Processing Letters, 14(3):124– 127, May 16, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elgot:1982:SSA

[EPS82] Calvin C. Elgot, Alan J. Perlis, and Lawrence Snyder. A syntax-free semantics for the APL operators. Information Processing Letters, 14(3): 128–131, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1973:LTS

[ER73] A. Ehrenfeucht and G. Rozenberg.

A limit theorem for sets of subwords in deterministic T0L languages. Information Processing Letters, 2(3): 70–73, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1979:FHB

[ER79a] Andrzej Ehrenfeucht and Grzegorz Rozenberg. Finding a homomorphism between two words is NPcomplete. Information Processing Letters, 9(2):86–88, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1979:OSG

[ER79b] Andrzej Ehrenfeucht and Grzegorz Rozenberg. An observation on scattered grammars. Information Processing Letters, 9(2):84–85, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1980:EIT

[ER80] A. Ehrenfeucht and G. Rozenberg. On the emptiness of the intersection of two D0S languages problem. Information Processing Letters, 10(4– 5):223–225, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erkio:1980:IMS

[Erk80] H. Erkio. Internal merge sorting with delayed selection. Information Processing Letters, 11(3):137– 140, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ershov:1977:PCP

[Ers77] A. P. Ershov. On the partial computation principle. Information Processing Letters, 6(2):38–41, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:1976:CT

[ES76] J. Engelfriet and S. Skyum. Copying theorems. Information Processing Letters, 4(6):157–161, March ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Endres:1977:FOP

[ES77] A. Endres and H.-G. Stork. FIFOoptimal placement on pages of independently referenced sectors. Information Processing Letters, 6(2): 46–49, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

FarinasDelCerro:1982:SDM

[Far82] L. Fariñas Del Cerro. A simple deduction method for modal logic. Information Processing Letters, 14(2): 49–51, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feyock:1976:NTE

[Fey76] Stefan Feyock. Noiselike transforms of ω-events. Information Processing Letters, 4(4):105–108, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friedman:1978:ESC

[FG78] E. P. Friedman and S. A. Greibach.

On equivalence and subclass containment problems for deterministic context-free languages. Information Processing Letters, 7(6):287– 290, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ferrari:1979:DLT

[FG79] M. Ferrari and G. Guida. DB: A LISP-type data base system. Information Processing Letters, 9(3): 126–134, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fortune:1979:NRN

[FH79] Steve Fortune and John Hopcroft.

A note on Rabin’s nearest-neighbor algorithm. Information Processing Letters, 8(1):20–23, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fridman:1979:ALS

[FHMS79] Felix J. Fridman, Glenn H. Holloway, Naftaly H. Minsky, and Josef Stein. Abstract FOR-loops over several aggregates. Information Processing Letters, 8(4):204–206, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Findler:1972:SNH

[Fin72] Nicholas V. Findler. Short note on a heuristic search strategy in longterm memory networks. Information Processing Letters, 1(5):191– 196, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fisher:1974:BWG

[Fis74] David A. Fisher. Bounded workspace garbage collection in an address-order preserving list processing environment. Information Processing Letters, 3(1):29–32, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fournier:1979:CAN

[FK79] Alain Fournier and Zvi Kedem. Comments on the all nearestneighbor problem for convex polygons. Information Processing Letters, 9(3):105–107, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LP78, YL79].

Fischer:1982:LBT

[FL82] Michael J. Fischer and Nancy A. Lynch. A lower bound for the time to assure interactive consistency. Information Processing Letters, 14(4): 183–186, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fournier:1979:CCH

[Fou79] Alain Fournier. Comments on convex hull of a finite set of points in two dimensions [Inform. Process. Lett. 7 (1978), no. 6, 296–298; MR 80b:68041]. Information Processing Letters, 8(4):173, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Byk78, DS79, Ov80].

Francez:1977:CFP

[Fra77] N. Francez. A case for a forward predicate transformer. Information Processing Letters, 6(6):196– 198, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraenkel:1979:PSL

[Fra79a] Aviezri S. Fraenkel. Paired sequential lists in a memory interval. Information Processing Letters, 8(1): 9–10, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franklin:1979:PLS

[Fra79b] W. Randolph Franklin. Padded lists: set operations in expected θ(log log N) time. Information Processing Letters, 9(4):161–166, November 20, 1979. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frederickson:1980:PAS

[Fre80] Greg N. Frederickson. Probabilistic analysis for simple one- and twodimensional bin packing algorithms. Information Processing Letters, 11 (4–5):156–161, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Floyd:1973:LTT

[FS73a] Robert W. Floyd and Alan J. Smith. A linear time two tape merge. Information Processing Letters, 2(5): 123–126 (or 123–125??), December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Francez:1973:NCC

[FS73b] Nissim Francez and Giora Slutzky. On the non-compactness of the class of program schemas. Information Processing Letters, 2(5):141– 142, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1979:IED

[FTM79] C. N. Fischer, K. C. Tai, and D. R. Milton. Immediate error detection in strong LL(1) parsers. Information Processing Letters, 8(5): 261–266, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friedman:1976:GCH

[FW76a] Daniel P. Friedman and David S. Wise. Garbage collecting a heap which includes a scatter table. Information Processing Letters, 5(6):161– 164, December ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [FW77].

Friedman:1976:ODI

[FW76b] Daniel P. Friedman and David S.

Wise. Output driven interpretation of recursive programs, or writing creates and destroys data structures. Information Processing Letters, 5(6):155–160, December ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [FW79a].

Friedman:1977:EGC

[FW77] Daniel P. Friedman and David S.

Wise. Erratum: “Garbage Collecting a Heap Which Includes a Scatter Table”. Information Processing Letters, 6(2):72, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [FW76a].

Friedman:1979:EOD

[FW79a] Daniel P. Friedman and David S.

Wise. Erratum: “Output Driven Interpretation of Recursive Programs, or Writing Creates and Destroys Data Structures”. Information Processing Letters, 9(2):101, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [FW76b].

Friedman:1979:RCC

[FW79b] Daniel P. Friedman and David S.

Wise. Reference counting can manage the circular environments of mutual recursion. Information Processing Letters, 8(1):41–45, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franchi-Zannettacci:1982:ETS

[FZ82] Paul Franchi-Zannettacci. An extension to trees of the Sardinas and Patterson algorithm. Information Processing Letters, 14(4):168–173, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabow:1976:NDC

[Gab76a] H. N. Gabow. A note on degreeconstrained star subgraphs of bipartite graphs. Information Processing Letters, 5(6):165–167, December ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabow:1976:SIB

[Gab76b] H. N. Gabow. Some improved bounds on the number of 1-factors of n-connected graphs. Information Processing Letters, 5(4):113– 115, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galil:1976:TFS

[Gal76] Zvi Galil. Two fast simulations which imply some fast string matching and palindrome-recognition algorithms. Information Processing Letters, 4(4):85–87, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gavril:1975:ATC

[Gav75] Fanica Gavril. An algorithm for testing chordality of graphs. Information Processing Letters, 3(4): 110–112, March ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gavril:1977:TEB

[Gav77] Fanica Gavril. Testing for equality between maximum matching and minimum node covering. Information Processing Letters, 6(6):199– 202, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gelenbe:1971:TTR

[Gel71] E. Gelenbe. The two-thirds rule for dynamic storage allocation under equilibrium. Information Processing Letters, 1(2):59–60, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gelenbe:1982:SP

[GG82] E. Gelenbe and D. Gardy. On the size of projections. Information Processing Letters, 14(1):18–21, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghezzi:1975:LGS

[Ghe75] C. Ghezzi. LL(1) grammars supporting an efficient error handling. Information Processing Letters, 3 (6):174–176, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gurari:1979:SCR

[GI79] Eitan M. Gurari and Oscar H.

Ibarra. On the space complexity of recursive algorithms. Information Processing Letters, 8(5):267– 271, June 11, 1979. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goto:1977:PHA

[GIG77] E. Goto, T. Ida, and T. Gunji. Parallel hashing algorithms. Information Processing Letters, 6(1):8– 13, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gini:1982:ASI

[Gin82] Guiseppina Gini. The automatic synthesis of iterative programs. Information Processing Letters, 14(2): 67–73, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garey:1978:TSP

[GJPT78] M. R. Garey, D. S. Johnson, F. P. Preparata, and R. E. Tarjan. Triangulating a simple polygon. Information Processing Letters, 7(4): 175–179, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galand:1975:SIS

[GL75] S. Galand and C. Loncour. Structured implementation of symbolic execution: a first part in a program verifier. Information Processing Letters, 3(4):97–103, March ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gries:1980:CFN

[GL80] David Gries and Gary Levin. Computing Fibonacci numbers (and similarly defined functions) in log time. Information Processing Letters, 11 (2):68–69, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gunn:1979:IC

[GM79] Hamish I. E. Gunn and Ronald Morrison. On the implementation of constants. Information Processing Letters, 9(1):1–4, July 20, 1979. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Graham:1982:SER

[GM82] M. H. Graham and A. O. Mendelzon. Strong equivalence of relational expressions under dependencies. Information Processing Letters, 14(2): 57–62, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Germano:1975:SSR

[GMS75] G. Germano and A. Maggiolo-Schettini. Sequence-to-sequence recursiveness. Information Processing Letters, 4(1):1–6, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonnet:1978:NDA

[Gon78] G. H. Gonnet. Notes on the derivation of asymptotic expressions from summations. Information Processing Letters, 7(4):165–169, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Gon79].

Gonnet:1979:END

[Gon79] G. H. Gonnet. Erratum to “Notes on the derivation of asymptotic expressions from summations” [Inform. Process. Lett. 7(4), June 1978, pp.

165–169]. Information Processing Letters, 8(1):54, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Gon78].

Gairola:1976:DIS

[GR76] B. K. Gairola and V. Rajaraman. A distributed index sequential access method. Information Processing Letters, 5(1):1–5, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonnet:1977:ISS

[GR77] G. H. Gonnet and L. D. Rogers. The interpolation-sequential search algorithm. Information Processing Letters, 6(4):136–139, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Graham:1972:EAD

[Gra72] R. L. Graham. An efficient algorithm for determining the convex hull of a finite planar set. Information Processing Letters, 1(4): 132–133, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [GS87].

Grant:1977:NVR

[Gra77a] J. Grant. Null values in a relational data base. Information Processing Letters, 6(5):156–157, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grant:1977:ARE

[Gra77b] P. W. Grant. Addendum: “Recognition of E0L languages in less than quatric time”. Information Processing Letters, 6(6):228, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Gra77c].

Grant:1977:REL

[Gra77c] P. W. Grant. Recognition of E0L languages in less than quatric time. Information Processing Letters, 6 (5):174–176 (or 174–175??), October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See addendum [Gra77b].

Grabowski:1979:USP

[Gra79a] Jan Grabowski. The unsolvability of some Petri net language problems. Information Processing Letters, 9(2):60–63, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grant:1979:PVT

[Gra79b] John Grant. Partial values in a tabular database model. Information Processing Letters, 9(2):97–99, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grabowski:1980:DPV

[Gra80] Jan Grabowski. The decidability of persistence for vector addition systems. Information Processing Letters, 11(1):20–23, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gries:1972:PI

[Gri72] D. Gries. Programming by induction. Information Processing Letters, 1(3):100–107, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guida:1976:SPR

[GS76] Giovanni Guida and Marco Somalvico. Semantics in problem representation and search. Information Processing Letters, 5(5):141– 145, November ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gries:1987:NGC

[GS87] David Gries and Ivan Stojmenovic. A note on Graham’s convex hull algorithm. Information Processing Letters, 25(5):323–327, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Gra72].

Garey:1978:LTA

[GT78] Michael R. Garey and Robert E. Tarjan. A linear-time algorithm for finding all feedback vertices. Information Processing Letters, 7(6): 274–276, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Glushkov:1978:CPT

[GTY78] V. M. Glushkov, G. E. Tseytlin, and E. L. Yushchenko. Certain problems of the theory of structured programs schemes. Information Processing Letters, 7(6):256–260, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gunn:1982:CTT

[Gun82] Hamish I. E. Gunn. Compile time type checking of structure field accessing. Information Processing Letters, 14(1):22–25, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giegerich:1978:COP

[GW78] Robert Giegerich and Reinhard Wilhelm. Counter-one-pass features in one-pass compilation: a formalization using attribute grammars. Information Processing Letters, 7(6): 279–284, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hansal:1974:SDP

[Han74] A. Hansal. ‘Software devices’ for processing graphs using PL/1 compile time facilities. Information Processing Letters, 2(6):171–179, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hansen:1975:UTC

[Han75] Per Brinch Hansen. Universal types in Concurrent Pascal. Information Processing Letters, 3(6):165– 166, July ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hansen:1976:CAL

[Han76a] P. Hansen. A cascade algorithm for the logical closure of a set of binary relations. Information Processing Letters, 5(2):50–54, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Han76b].

Hansen:1976:ECA

[Han76b] P. Hansen. Erratum: “A cascade algorithm for the logical closure of a set of binary relations”. Information Processing Letters, 5(4):124, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Han76a].

Hanson:1976:SVB

[Han76c] D. R. Hanson. A simple variant of the boundary-tag algorithm for the allocation of coroutine environments. Information Processing Letters, 4(4):109–112 (or 109–111??), January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hansen:1978:PAO

[Han78] W. J. Hansen. A predecessor algorithm for ordered lists. Information Processing Letters, 7(3):137– 138, April 28, 1978. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanson:1980:CIL

[Han80] D. R. Hanson. Code improvement via lazy evaluation. Information Processing Letters, 11(4–5):163–167, December 12, 1980. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harel:1979:ETR

[Har79a] D. Harel. Erratum: “Two results on process logic”. Information Processing Letters, 9(3):160, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Har79b].

Harel:1979:TRP

[Har79b] David Harel. Two results on process logic. Information Processing Letters, 8(4):195–198, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [Har79a].

Herschberg:1977:CFA

[HB77] I. S. Herschberg and J. C. A.

Boekhorst. Concurrent file access under unpredictability. Information Processing Letters, 6(6):203– 208, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hellman:1981:ACA

[Hel81] M. E. Hellman. Another cryptanalytic attack on “A cryptosystem for multiple communication” [Inform. Process. Lett. 10(4–5), 5 July 1980, pp. 180–183]. Information Processing Letters, 12(4):182–183, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LM80b, Mei81].

Henhapl:1973:TMG

[Hen73] W. Henhapl. A transformation of marked graphs. Information Processing Letters, 2(1):26–29, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Henno:1979:DMF

[Hen79] J. Henno. The depth of monotone functions in multivalued logic. Information Processing Letters, 8(4): 176–177, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:1973:UCC

[Her73] Gabor T. Herman. On universal computer-constructors. Information Processing Letters, 2(3):61–64, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hikita:1977:AFP

[HG77] T. Hikita and E. Goto. An O(N) algorithm for finding periodicity of a sequence using hash coding. Information Processing Letters, 6(2): 69–71, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirschberg:1978:ITL

[Hir78] D. S. Hirschberg. An informationtheoretic lower bound for the longest common subsequence problem. Information Processing Letters, 7(1): 40–41, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Henhapl:1971:RTM

[HJ71] W. Henhapl and C. B. Jones. A runtime mechanism for referencing variables. Information Processing Letters, 1(1):14–16, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huits:1979:PSD

[HK79] Martin Huits and Vipin Kumar. The practical significance of distributive partitioning sort. Information Processing Letters, 8(4):168–169, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See original paper [Dob78b] and rebuttal [Dob79], and further debate [Bur78, Dob78a].

Heuer:1973:NSB

[HM73] H. Heuer and I. N. Molchanov. Numerical solution of a boundary problem for equations of elastic equilibrium of bodies in transferences. Information Processing Letters, 2(5): 148–151, December ??, 1973. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hitotumatu:1979:TIB

[HN79] Hirosi Hitotumatu and Kohei Noshita. A technique for implementing backtrack algorithms and its application. Information Processing Letters, 8(4):174–175, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoare:1973:GCL

[Hoa73] C. A. R. Hoare. A general conservation law for queueing disciplines. Information Processing Letters, 2(3): 82–85, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoare:1974:OSS

[Hoa74] C. A. R. Hoare. Optimization of store size for garbage collection. Information Processing Letters, 2(6): 165–166, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Horowitz:1972:EFM

[Hor72a] E. Horowitz. Erratum: “A fast method for interpolation using preconditioning”. Information Processing Letters, 1(5):216, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Hor72b].

Horowitz:1972:FMI

[Hor72b] E. Horowitz. A fast method for interpolation using preconditioning. Information Processing Letters, 1 (4):157–163, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Hor72a].

Hosono:1975:CSL

[Hos75] Chiharu Hosono. On the cardinality of some lattices. Information Processing Letters, 3(3):67–68, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Howden:1978:LGS

[How78] W. E. Howden. Lindenmayer grammars and symbolic testing. Information Processing Letters, 7(1):36–39, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hyafil:1976:COB

[HR76] L. Hyafil and R. L. Rivest. Constructing optimal binary decision trees is NP-complete. Information Processing Letters, 5(1):15–17, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Holnapy:1973:GMF

[HSB73] D. Holnapy, M. Szöts, and A. Botár. A generalization of the method of finite differences. Information Processing Letters, 2(2):55–59, June ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hopcroft:1971:ADI

[HT71] J. Hopcroft and R. E. Tarjan. A V^{2} algorithm for determining isomorphism of planar graphs. Information Processing Letters, 1(1):32– 34, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huet:1978:AGB

[Hue78] G. Huet. An algorithm to generate the basis of solutions to homogeneous linear Diophantine equations. Information Processing Letters, 7(3):144–147 (or 144–146??), April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hughes:1978:EVA

[Hug78] Charles E. Hughes. The equivalence of vector addition systems to a subclass of Post canonical systems. Information Processing Letters, 7(4): 201–204, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hyafil:1975:ACS

[HV75] L. Hyafil and J. P. Van de Wiele. On the additive complexity of specific polynomials. Information Processing Letters, 4(2):45–47, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hennell:1976:PA

[HWH76] M. A. Hennell, M. R. Woodward, and D. Hedley. On program analysis. Information Processing Letters, 5(5):136–140, November ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1974:NSS

[Iba74] Oscar H. Ibarra. A note on semilinear sets and bounded-reversal multihead pushdown automata. Information Processing Letters, 3(1):25–28, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Illies:1978:CGA

[Ill78] N. Illies. A counterexample to the generalized Aanderaa-Rosenberg conjecture. Information Processing Letters, 7(3):154–155, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1980:NPC

[IMR80] Oscar H. Ibarra, Shlomo Moran, and Louis E. Rosier. A note on the parallel complexity of computing the rank of order n matrices. Information Processing Letters, 11(4–5):162, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Istrail:1978:TSG

[Ist78] S. Istrail. Tag systems generating Thue irreducible sequences. Information Processing Letters, 7(3): 129–131, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Innes:1976:IAP

[IT76] Donald R. Innes and Shalom Tsur. Interval analysis, pagination and program locality. Information Processing Letters, 5(4):91–96, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1979:NBP

[IT79a] Katsushi Inoue and Itsuo Takanami.

A note on bottom-up pyramid acceptors. Information Processing Letters, 8(1):34–37, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1979:NCC

[IT79b] Katsushi Inoue and Itsuo Takanami.

A note on cyclic closure operations. Information Processing Letters, 8 (1):15–16, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1980:NDP

[IT80] Katsushi Inoue and Itsuo Takanami.

A note on decision problems for three-way two-dimensional finite automata. Information Processing Letters, 10(4–5):245–248, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itai:1979:RAC

[Ita79a] Alon Itai. A randomized algorithm for checking equivalence of circular lists. Information Processing Letters, 9(3):118–121, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itano:1979:RPS

[Ita79b] K. Itano. Reduction of page swaps on the two dimensional transforms in a paging environment. Information Processing Letters, 9(3):137– 140, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1978:NTF

[ITN78] Katsushi Inoue, Itsuo Takanami, and Akira Nakamura. A note on two-dimensional finite automata. Information Processing Letters, 7 (1):49–52, January 12, 1978. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jaffe:1978:CPC

[Jaf78] J. Jaffe. Counting productions in context-free derivations. Information Processing Letters, 7(4):196– 200, June ??, 1978. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jakobsson:1978:HCB

[Jak78] M. Jakobsson. Huffman coding in bit-vector compression. Information Processing Letters, 7(6):304– 307, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jakobsson:1982:EHB

[Jak82] M. Jakobsson. Evaluation of a hierarchical bit-vector compression technique. Information Processing Letters, 14(4):147–149, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jantzen:1979:NVG

[Jan79] M. Jantzen. A note on vector grammars. Information Processing Letters, 8(1):32–33, January 2, 1979.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jarvis:1973:ICH

[Jar73] R. A. Jarvis. On the identification of the convex hull of a finite set of points in the plane. Information Processing Letters, 2(1):18–21, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Javor:1977:ATA

[Jav77] A. Javor. An adaptive time advancement algorithm for discrete simulation. Information Processing Letters, 6(3):83–86, April ??, 1977. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jackson:1974:IPM

[JD74] L. Wayne Jackson and Subrata Dasgupta. The identification of parallel micro-operations. Information Processing Letters, 2(6):180–184, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jou:1982:CRR

[JF82] Jiann H. Jou and Patrick C. Fischer.

The complexity of recognizing 3NF relation schemes. Information Processing Letters, 14(4):187–190, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jarzabek:1975:LRG

[JK75] Stanislaw Jarzabek and Tomasz Krawczyk. LL-regular grammars. Information Processing Letters, 4 (2):31–37, November ??, 1975. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jackowski:1979:CSD

[JKS79] Boguslaw L. Jackowski, Ryszard Kubiak, and Stefan Sokołowski. Complexity of sorting by distributive partitioning. Information Processing Letters, 9(2):100, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:1975:PQU

[Joh75] Donald B. Johnson. Priority queues with update and finding minimum spanning trees. Information Processing Letters, 4(3):53–57, December ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:1980:CVA

[Joh80] Robert B. Johnson, Jr. The complexity of a VLSI adder. Information Processing Letters, 11(2):92–93, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jones:1977:NLT

[Jon77] N. D. Jones. A note on linear time simulation of deterministic two-way pushdown automata. Information Processing Letters, 6(4): 110–112, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jonkers:1979:FGC

[Jon79] H. B. M. Jonkers. A fast garbage compaction algorithm. Information Processing Letters, 9(1):26–30, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jovanovich:1974:NMF

[Jov74] A. D. Jovanovich. Note on a modification of the fundamental cycles finding algorithm. Information Processing Letters, 3(1):33–??, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jammel:1980:ECD

[JS80] Alfons J. Jammel and Helmut G.

Stiegler. On expected costs of deadlock detection. Information Processing Letters, 11(4–5):229–231, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kang:1975:SPC

[KA75] A. N. C. Kang and D. A. Ault.

Some properties of a centroid of a free tree. Information Processing Letters, 4(1):18–20, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kameda:1975:VRR

[Kam75] T. Kameda. On the vector representation of the reachability in planar directed graphs. Information Processing Letters, 3(3):75–77, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kanellakis:1980:CCC

[Kan80] Paris C. Kanellakis. On the computational complexity of cardinality constraints in relational databases. Information Processing Letters, 11 (2):98–101, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karlgren:1971:SRS

[Kar71] H. Karlgren. Stacking without really stacking when reducing categorical expressions. Information Processing Letters, 1(1):21–22, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karhumaki:1973:EPS

[Kar73] Juhani Karhumäki. An example of a PD2L-system with the growth type 21/2. Information Processing Letters, 2(5):131–134, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kasyanov:1973:SPF

[Kas73] V. N. Kasyanov. Some properties of fully reducible graphs. Information Processing Letters, 2(4):113– 117, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kawai:1980:FSP

[Kaw80] Hiroya Kawai. A formal system for parallel programs in discrete time and space. Information Processing Letters, 11(4–5):204–210, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaye:1976:GCS

[Kay76] R. Kaye. A Gray Code for set partitions. Information Processing Letters, 5(6):171–173, December ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kral:1973:NNS

[KD73] J. Král and J. Demner. A note on the number of states of the De Remer’s recognizer. Information Processing Letters, 2(1):22–23, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keir:1975:SSR

[Kei75] R. A. Keir. Should the stable rounding rule be radix-dependent? Information Processing Letters, 3(6): 188–189, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knuth:1971:ENA

[KF71a] Donald E. Knuth and R. W. Floyd.

Erratum: “Notes on avoiding ‘go to’ statements”. Information Processing Letters, 1(1):23–31, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [KF71b].

Knuth:1971:NAG

[KF71b] Donald E. Knuth and R. W. Floyd.

Notes on avoiding “go to” statements. Information Processing Letters, 1(1):23–31, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [KF71a]. Reprinted in Writings of the Revolution, E. Yourdon, ed. (New York: Yourdon Press, 1982), 153–162.

Kambayashi:1978:LSS

[KHTY78] Yahiko Kambayashi, Takaki Hayashi, Yoshikazu Tanaka, and Shuzo Yajima. A linear storage space algorithm for a reference structure index. Information Processing Letters, 7(2):66–71, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kinber:1980:IPD

[Kin80] E. B. Kinber. On inclusion problem for deterministic multitape automata. Information Processing Letters, 11(3):144–146, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kittler:1975:DJD

[Kit75] J. Kittler. On the divergence and the Joshi dependence measure in feature selection. Information Processing Letters, 3(5):135–137, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kittler:1977:MDC

[Kit77] J. Kittler. A method for determining class subspaces. Information Processing Letters, 6(3):77–79, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koplowitz:1978:MEC

[KJ78] J. Koplowitz and D. Jouppi. A more efficient convex hull algorithm. Information Processing Letters, 7(1): 56–57, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knuth:1972:EOS

[KK72] Donald E. Knuth and E. B. Kaehler. An experiment in optimal sorting. Information Processing Letters, 1 (4):173–176, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krzemien:1976:AGL

[KŁ76] Roman Krzemień and Andrezej Łukasiewicz. Automatic generation of lexical analyzers in a compilercompiler. Information Processing Letters, 4(6):165–168, March ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karayiannis:1978:CDC

[KL78] Andreas Karayiannis and Georghios Loizou. Cycle detection in critical path networks. Information Processing Letters, 7(1):15–19, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kessels:1979:TIC

[KM79] J. L. W. Kessels and A. J. Martin.

Two implementations of the conditional critical region using a split binary semaphore. Information Processing Letters, 8(2):67–71, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kandzia:1980:CBC

[KM80] Peter Kandzia and Margret Mangelmann. On covering Boyce-Codd normal forms. Information Processing Letters, 11(4–5):218–223, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knuth:1977:GDA

[Knu77] Donald E. Knuth. A generalization of Dijkstra’s algorithm. Information Processing Letters, 6(1):1–5, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming. CSLI Lecture Notes Number 27. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xvi + 368 pp. LCCN QA76.6 .K644 1992. US$24.95 (paperback), US$59.95 (hardcover). Distributed by the University of Chicago Press.

Ko:1982:SOP

[Ko82] Ker-I Ko. Some observations on the probabilistic algorithms and NPhard problems. Information Processing Letters, 14(1):39–43, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobayashi:1974:NEE

[Kob74] Kojiro Kobayashi. A note on extending equivalence theories of algorithms. Information Processing Letters, 3(2):54–56, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobuchi:1977:TCT

[Kob77] Y. Kobuchi. Two characterization theorems of locally catenative developmental systems. Information Processing Letters, 6(4):120–124, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1982:GNO

[Kor82] J. F. Korsh. Growing nearly optimal binary search trees. Information Processing Letters, 14(3):139– 143, May 16, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krepyshev:1972:SDO

[KP72] G. B. Krepyshev and Ya. A. Pollack.

Synthesis of a discrete-time optimal filter algorithm with reduced sensitivity to deviations of a priori statistics. Information Processing Letters, 1(5):211–215, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krishnamurthy:1976:NPF

[KR76] M. S. Krishnamurthy and H. R.

Ramesha Chandra. A note on precedence functions. Information Processing Letters, 4(4):99–100, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kral:1972:ENA

[Krá72a] J. Král. Erratum: “A new additive pseudorandom number generator for extremely short wordlength”. Information Processing Letters, 1(5):216, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Krá72b].

Kral:1972:NAP

[Krá72b] J. Král. A new additive pseudorandom number generator for extremely short word-length. Information Processing Letters, 1(4):164–167, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Krá72a].

Krawczyk:1980:ECM

[Kra80] Tomasz Krawczyk. Error correction by mutational grammars. Information Processing Letters, 11(1):9–15, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krichevsky:1982:LEF

[Kri82] R. E. Krichevsky. Letter to the editor: “A family of codes commutatively equivalent to prefix codes” [Inform. Process. Lett. 12 (1981), no. 1, 1–4; MR 82j:94021] by S. Mauceri and A. Restivo. Information Processing Letters, 14(5):238, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [MR81].

Krogdahl:1973:DSA

[Kro73] Stein Krogdahl. A dynamic storage allocation problem. Information Processing Letters, 2(4):96–99, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kruijer:1979:SSS

[Kru79] H. S. M. Kruijer. Self-stabilization (in spite of distributed control) in tree-structured systems. Information Processing Letters, 8(2):91–95, February 15, 1979. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knuth:1974:ESP

[KS74a] Donald E. Knuth and Jayme L.

Szwarcfiter. Erratum: “A Structured Program to Generate All Topological Sorting Arrangements”. Information Processing Letters, 3 (2):64, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [KS74b].

Knuth:1974:SPG

[KS74b] Donald E. Knuth and Jayme L.

Szwarcfiter. A structured program to generate all topological sorting arrangements. Information Processing Letters, 2(6):153–157, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [KS74a]. Reprinted with revisions in [Knu92, pp. 91–97].

King:1982:OAS

[KST82] K. N. King and Barbara Smith-Thomas. An optimal algorithm for sink-finding. Information Processing Letters, 14(3):109–111, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucera:1982:PCC

[Kuč82] Luděk Kučera. Parallel computation and conflicts in memory access. Information Processing Letters, 14 (2):93–96, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum and addendum [Kuč83].

Kucera:1983:EAP

[Kuč83] Luděk Kučera. Erratum and addendum to: “Parallel computation and conflicts in memory access” [Inform. Process. Lett. 14 (1982), no. 2, 93–96; MR 83g:68038]. Information Processing Letters, 17(2):107, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Kuč82].

Kung:1973:NUB

[Kun73] H. T. Kung. A new upper bound on the complexity of derivative evaluation. Information Processing Letters, 2(5):146–147, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kundu:1976:LAH

[Kun76] S. Kundu. A linear algorithm for the Hamiltonian completion number of a tree. Information Processing Letters, 5(2):55–57, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kundu:1977:STN

[Kun77] S. Kundu. Sorting tree, nestling tree and inverse permutation. Information Processing Letters, 6(3):94–96, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kundu:1979:IVT

[Kun79] Sukhamay Kundu. An intermediatevalue theorem for optimum tree valuation. Information Processing Letters, 8(3):141–145, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lamport:1976:CSA

[Lam76] L. Lamport. Comments on “A synchronization anomaly”. Information Processing Letters, 4(4):88–89, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LT75].

Lao:1979:NDS

[Lao79] Marek J. Lao. A new data structure for the UNION-FIND problem. Information Processing Letters, 9(1):39–45, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laut:1980:SPI

[Lau80] A. Laut. Safe procedural implementations of algebraic types. Information Processing Letters, 11(4–5): 147–151, December 12, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lawler:1976:NCC

[Law76] E. L. Lawler. A note on the complexity of the chromatic number problem. Information Processing Letters, 5(3):66–67, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lazos:1975:CSR

[Laz75] Constantine Lazos. A comparison of simulation results and a mathematical model of a multiprogramming system. Information Processing Letters, 3(5):127–134, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucena:1976:TSE

[LC76] C. J. Lucena and D. D. Cowan.

Toward a system’s environment for computer assisted programming. Information Processing Letters, 5(2):

35–40, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1980:ATL

[LD80] Lishing Liu and Alan Demers. An algorithm for testing lossless join property in relational databases. Information Processing Letters, 11(2): 73–76, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leipala:1979:GBS

[Lei79] Timo Leipälä. On a generalization of binary search. Information Processing Letters, 8(5):230–233, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leiss:1980:NSS

[Lei80] E. Leiss. A note on a signature system based on probabilistic logic. Information Processing Letters, 11(2): 110–113, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lenstra:1979:MPT

[Len79] H. W. Lenstra, Jr. Miller’s primality test. Information Processing Letters, 8(2):86–88, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1979:BLS

[Leu79] Joseph Y.-T.-T. Leung. Bounds on list scheduling of UET tasks with restricted resource constraints. Information Processing Letters, 9(4):167– 170, November 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levy:1975:COP

[Lev75] M. R. Levy. Complete operator precedence. Information Processing Letters, 4(2):38–40, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lang:1976:SUL

[LF76] Tomas Lang and Eduardo B. Fernandez. Scheduling of unit-length independent tasks with execution constraints. Information Processing Letters, 4(4):95–98, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lindstrom:1973:SLS

[Lin73] Gary Lindstrom. Scanning list structures without stacks or tag bits. Information Processing Letters, 2 (2):47–51, June ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Linna:1976:DNC

[Lin76] M. Linna. The D0L-ness for contextfree languages is decidable. Information Processing Letters, 5(5):149– 151, November ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipski:1977:OMP

[Lip77] W. Lipski, Jr. One more polynomial complete consecutive retrieval problem. Information Processing Letters, 6(3):91–93, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lloyd:1980:ELS

[Llo80a] E. L. Lloyd. Errata: “List scheduling bounds for UET systems with resources” [Inform. Process. Lett. 10 (1980), no. 1, 28–31; MR 81a:68042]. Information Processing Letters, 11 (1):57, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Llo80b].

Lloyd:1980:LSB

[Llo80b] Errol L. Lloyd. List scheduling bounds for UET systems with resources. Information Processing Letters, 10(1):28–31, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also errata [Llo80a].

Langmaack:1979:FTP

[LLW79] Hans Langmaack, Wolfram Lippe, and Franz Wagner. The formal termination problem for programs with finite ALGOL 68-modes. Information Processing Letters, 9(3): 155–159, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipton:1978:BMC

[LM78] Richard J. Lipton and Raymond E. Miller. A batching method for coloring planar graphs. Information Processing Letters, 7(4):185–188, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1980:NPS

[LM80a] Joseph Y.-T. Leung and M. L. Merrill. A note on preemptive scheduling of periodic, real-time tasks. Information Processing Letters, 11(3):

115–118, November 18, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1980:CMC

[LM80b] F. Luccio and S. Mazzone. A cryptosystem for multiple communication. Information Processing Letters, 10(4–5):180–183, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See notes [Mei81, Hel81].

Lassez:1982:FPT

[LNS82] J.-L. Lassez, V. L. Nguyen, and E. A. Sonenberg. Fixed point theorems and semantics: a folk tale. Information Processing Letters, 14(3): 112–116, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loizou:1979:MSD

[Loi79] Georghios Loizou. Mathematical solution for a data processing system. Information Processing Letters, 8(5):252–256, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loizou:1980:CFA

[Loi80] G. Loizou. On a cycle finding algorithm. Information Processing Letters, 11(1):33–36, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loui:1980:NPG

[Lou80] Michael C. Loui. A note on the pebble game. Information Processing Letters, 11(1):24–26, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lozinskii:1976:PSO

[Loz76] E. L. Lozinskii. On a problem in storage optimization. Information Processing Letters, 5(3):78–80, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1976:SCR

[LP76] F. Luccio and F. P. Preparata. Storage for consecutive retrieval. Information Processing Letters, 5(3): 68–71, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1978:ANN

[LP78] D. T. Lee and F. P. Preparata. The all nearest-neighbor problem for convex polygons. Information Processing Letters, 7(4):189–192, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [YL79, FK79].

Lee:1974:LSD

[LR74] K. P. Lee and G. Rozenberg. The length sets of D0L languages are uniformly bounded. Information Processing Letters, 2(6):185–188, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luque:1980:TAM

[LR80] E. Luque and A. Ripoll. Tuning architecture via microprogramming. Information Processing Letters, 11 (2):102–109, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lagarias:1982:MSR

[LS82] J. C. Lagarias and D. E. Swartwout.

Minimal storage representations for binary relations. Information Processing Letters, 14(2):63–66, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipton:1975:SA

[LT75] Richard J. Lipton and Robert W.

Tuttle. A synchronization anomaly. Information Processing Letters, 3 (3):65–66, January ??, 1975. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [Lam76].

Lukaszewicz:1977:FG

[Łuk77] L. Łukaszewicz. Functional grammars. Information Processing Letters, 6(5):146–150, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luk:1979:PMM

[Luk79] W. S. Luk. ‘Possible’ membership of a multivalued dependency in a relational database. Information Processing Letters, 9(2):80–83, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1979:LMP

[LY79] D. T. Lee and C. C. Yang. Location of multiple points in a planar subdivision. Information Processing Letters, 9(4):190–193, November 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lyon:1979:BSS

[Lyo79] Gordon Lyon. Batch scheduling from short lists. Information Processing Letters, 8(2):57–59, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McCallum:1979:LAF

[MA79] Duncan McCallum and David Avis. A linear algorithm for finding the convex hull of a simple polygon. Information Processing Letters, 9(5): 201–206, December 16, 1979. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meijer:1980:DAN

[MA80] Henk Meijer and Selim G. Akl. The design and analysis of a new hybrid sorting algorithm. Information Processing Letters, 10(4–5):213–218, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madsen:1977:EFD

[Mad77] J. Madsen. An experiment in formal definition of operating system facilities. Information Processing Letters, 6(6):187–189, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makarov:1979:UDC

[Mak79] O. M. Makarov. Using duality to compute the pair of matrix products QY and Y^{T} Q over a commutative ring. Information Processing Letters, 8(1):46–49, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manacher:1976:APM

[Man76a] G. K. Manacher. An application of pattern matching to a problem in geometrical complexity. Information Processing Letters, 5(1):6–7, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mandrioli:1976:RCF

[Man76b] D. Mandrioli. n-reconstructability of context-free grammars. Information Processing Letters, 5(2):58–62, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marsaglia:1976:IFM

[MAP76] G. Marsaglia, K. Ananthanarayanan, and N. J. Paul. Improvements on fast methods for generating normal random variables. Information Processing Letters, 5(2):27–30, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matwin:1977:EIG

[Mat77a] S. Matwin. An experimental investigation of Geschke’s method of global program optimization. Information Processing Letters, 6(6):177– 179, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matwin:1977:CST

[Mat77b] S. Matwin. On the completeness of a set of transformations optimizing linear programs. Information Processing Letters, 6(5):165–167, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mathon:1979:NGI

[Mat79] Rudolf Mathon. A note on the graph isomorphism counting problem. Information Processing Letters, 8(3): 131–132, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maurer:1971:SPG

[Mau71] H. A. Maurer. The solution of a problem by Ginsburg. Information Processing Letters, 1(1):7–11 (or 7– 10??), February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mayoh:1972:RS

[May72] B. H. Mayoh. Recursion and stacks. Information Processing Letters, 1 (3):115–116, February ??, 1972. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majster-Cederbaum:1980:SRB

[MC80] Mila E. Majster-Cederbaum. A simple relation between relational and predicate transformer semantics for nondeterministic programs. Information Processing Letters, 11(4–5): 190–192, December 12, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McColl:1978:MDM

[McC78] W. F. McColl. The maximum depth of monotone formulae. Information Processing Letters, 7(2):65–??, February 28, 1978. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meertens:1976:SST

[Mee76] L. Meertens. A space-saving technique for assigning ALGOL 68 multiple values. Information Processing Letters, 5(4):97–99, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:1976:BLR

[Meh76] K. Mehlhorn. Bracket-languages are recognizable in logarithmic space. Information Processing Letters, 5 (6):168–170, December ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meijer:1981:NCM

[Mei81] H. Meijer. A note on “A cryptosystem for multiple communication” [Inform. Process. Lett. 10(4– 5), 5 July 1980, pp. 180–183]. Information Processing Letters, 12(4): 179–181, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LM80b, Hel81].

Merret:1971:GPM

[Mer71] T. H. Merret. General programs for management systems. Information Processing Letters, 1(1):17– 20, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Merret:1977:RPL

[Mer77] T. H. Merret. Relations as programming language elements. Information Processing Letters, 6(1): 29–33, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Merret:1978:QLS

[Mer78] T. H. Merret. QT logic: simpler and more expressive than predicate calculus. Information Processing Letters, 7(6):251–255, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Molchanov:1973:CDS

[MG73] I. N. Molchanov and E. F. Galba. On the convergence of difference schemes approximating a plane static problem of the theory of elasticity with mixed boundary conditions. Information Processing Letters, 2(4):118–122, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mitchell:1979:LAE

[MH79] Sandra L. Mitchell and Stephen Hedetniemi. Linear algorithms for edge-coloring trees and unicyclic graphs. Information Processing Letters, 9(3):110–112, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Molchanov:1973:OCI

[MI73] I. N. Molchanov and M. F. Iakovlev. On one class of iterative methods for obtaining the generalized solution of non-consistent systems of linear algebraic equations. Information Processing Letters, 2(3):86–90, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Middleton:1974:COP

[Mid74] A. G. Middleton. Cost-oriented program optimisation. Information Processing Letters, 2(6):167– 170, April ??, 1974. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minsky:1973:RBT

[Min73] Naftaly H. Minsky. Representation of binary trees on associative memories. Information Processing Letters, 2(1):1–5, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:1979:STT

[Mis79] Jayadev Misra. Space-time trade off in implementing certain set operations. Information Processing Letters, 8(2):81–85, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mitchell:1979:LAR

[Mit79] Sandra L. Mitchell. Linear algorithms to recognize outerplanar and maximal outerplanar graphs. Information Processing Letters, 9(5):229– 232, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Malhotra:1978:AFM

[MKM78] V. M. Malhotra, M. Pramodh Kumar, and S. N. Maheshwari. An O(V^{3}) algorithm for finding maximum flows in networks. Information Processing Letters, 7(6):277– 278, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Molchanov:1972:AIB

[MN72] I. N. Molchanov and L. D.

Nikolenko. On an approach to integrating boundary problems with a non-unique solution. Information Processing Letters, 1(4):168– 172, June ??, 1972. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moore:1979:MPT

[Moo79] J. Strother Moore. A mechanical proof of the termination of Takeuchi’s function. Information Processing Letters, 9(4):176–181, November 20, 1979. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morris:1979:SFS

[Mor79a] J. M. Morris. A starvation-free solution to the mutual exclusion problem. Information Processing Letters, 8(2):76–80, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [AJV86].

Morris:1979:TBT

[Mor79b] J. M. Morris. Traversing binary trees simply and cheaply. Information Processing Letters, 9(5):197– 200, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moran:1981:NSP

[Mor81] Shlomo Moran. A note on: “Shortest-path problem is not harder than matrix multiplication” [Inform. Process. Lett. 11 (1980), no. 3, 134–136; MR 81k:68036] by F. Romani. With a reply by Romani. Information Processing Letters, 13 (2):85–86, November 13, 1981. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Rom80].

Maurer:1976:IDU

[MOS76] H. A. Maurer, Th. Ottmann, and H.-W. W. Six. Implementing dictionaries using binary trees of very small height. Information Processing Letters, 5(1):11–14, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moura:1980:NGC

[Mou80] Arnaldo Moura. A note on grammatical covers. Information Processing Letters, 11(3):127–129, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Milgram:1971:NSC

[MR71] D. L. Milgram and A. Rosenfeld. A note on scattered context grammars. Information Processing Letters, 1 (2):47–50, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mauceri:1981:FCC

[MR81] S. Mauceri and A. Restivo. A family of codes commutatively equivalent to prefix codes. Information Processing Letters, 12(1):1–4, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See letter to the editor [Kri82].

Madsen:1976:MMD

[MRK76] Niel K. Madsen, Garry H. Rodrigue, and Jack I. Karush. Matrix multiplication by diagonals on a vector/ parallel processor. Information Processing Letters, 5(2):41–45, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Molchanov:1972:IMS

[MS72] I. N. Molchanov and N. I. Stepanets. Iterative methods for solving difference equations of the theory of elasticity not depending on the spacing of the difference net. Information Processing Letters, 1(5):205– 210, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:1975:OCP

[MT75] J. Misra and R. E. Tarjan. Optimal chain partitions of trees. Information Processing Letters, 4(1):24– 26, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mizumoto:1973:EFG

[MTT73] Masaharu Mizumoto, Junichi Toyoda, and Kohkichi Tanaka. Examples of formal grammars with weights. Information Processing Letters, 2(3):74–78, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muhlbacher:1979:FGG

[Müh79] Jörg Mühlbacher. F -factors of graphs: a generalized matching problem. Information Processing Letters, 8(4):207–214, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Munro:1971:EDT

[Mun71] I. Munro. Efficient determination of the transitive closure of a directed graph. Information Processing Letters, 1(2):56–58, July ??, 1971. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minker:1974:EAP

[MV74] Jack Minker and Gordon J. VanderBrug. The Earley algorithm as a problem representation. Information Processing Letters, 3(1):1–7, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manacher:1979:NGN

[MZ79] Glenn K. Manacher and Albert L.

Zobrist. Neither the Greedy nor the Delaunay triangulation of a planar point set approximates the optimal triangulation. Information Processing Letters, 9(1):31–34, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minker:1982:ELR

[MZ82] Jack Minker and Guy Zanon. An extension to linear resolution with selection function. Information Processing Letters, 14(4):191–194, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakamura:1980:RTD

[NI80] Akira Nakamura and Katsushi Inoue. A remark on two-dimensional finite automata: “Some properties of two-dimensional on-line tessellation acceptors” [Inform. Sci.

13 (1977), no. 2, 95–121; MR 80e:68143]. Information Processing Letters, 10(4–5):219–222, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nieminen:1975:HIT

[Nie75a] J. Nieminen. On homomorphic images of transition graphs. Information Processing Letters, 4(1):14– 15, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nieminen:1975:SOD

[Nie75b] J. Nieminen. Some observations on the determination of an upper bound for the clique number of a graph. Information Processing Letters, 4(1):16–17, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nijholt:1979:LRG

[Nij79] A. Nijholt. From left-regular to Greibach normal form grammars. Information Processing Letters, 9 (1):51–55, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nath:1982:PAC

[NM82] Dhruva Nath and S. N. Maheshwari. Parallel algorithms for the connected components and minimal spanning tree problems. Information Processing Letters, 14(1):7–11, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noshita:1978:EBD

[NMM78] Kohei Noshita, Etsuo Masuda, and Hajime Machida. On the expected behaviors of the Dijkstra’s shortest path algorithm for complete graphs. Information Processing Letters, 7 (5):237–243, August ??, 1978. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noga:1987:SPD

[Nog87] M. T. Noga. Sorting in parallel by double distributive partitioning. BIT, 27(3):340–348, September 1987. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=27&issue=3&spage=

340. See [Dob78b].

Noshita:1974:MSE

[Nos74] K. Noshita. Median selection of 9 elements in 14 comparisons. Information Processing Letters, 3(1):8– 12, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nozaki:1979:NCA

[Noz79] Akihiro Nozaki. A note on the complexity of approximative evaluation of polynomials. Information Processing Letters, 9(2):73–75, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nievergelt:1972:BWP

[NPWY72] J. Nievergelt, J. Pradels, C. K.

Wong, and P. C. Yue. Bounds on the weighted path length of binary trees. Information Processing Letters, 1(6):220–225, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Natarajan:1978:ODW

[NW78] K. S. Natarajan and Lee J. White. Optimum domination in weighted trees. Information Processing Letters, 7(6):261–265, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ogilvie:1972:UCS

[OO72] J. C. Ogilvie and C. L. Olson. On the use of complete subgraphs in cluster analysis. Information Processing Letters, 1(3):76–79, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Orenstein:1982:MTU

[Ore82] Jack A. Orenstein. Multidimensional tries used for associative searching. Information Processing Letters, 14(4):150–157, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Osborn:1979:TEC

[Osb79] Sylvia L. Osborn. Testing for existence of a covering Boyce-Codd normal form. Information Processing Letters, 8(1):11–14, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oulsnam:1979:CND

[Oul79] G. Oulsnam. Cyclomatic numbers do not measure complexity of unstructured programs. Information Processing Letters, 9(5):207– 211, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Overmars:1980:FCB

[Ov80] Mark H. Overmars and Jan van Leeuwen. Further comments on A. Bykat’s convex hull algorithm: “Convex hull of a finite set of points in two dimensions” [Inform. Process. Lett. 7 (1978), no. 6, 296–298; MR 80b:68041]. Information Processing Letters, 10(4–5):209–212, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Byk78, DS79, Fou79].

Pajunen:1980:TTL

[Paj80] Seppo Pajunen. On two theorems of Lenstra. Information Processing Letters, 11(4–5):224–228, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Papadimitriou:1979:ESR

[Pap79] Christos H. Papadimitriou. Efficient search for rationals. Information Processing Letters, 8(1):1–4, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paredaens:1978:EPR

[Par78] J. Paredaens. On the expressive power of the relational algebra. Information Processing Letters, 7(2): 107–111, February 28, 1978. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paun:1978:GCS

[Pău78] G. Păun. On the generative capacity of simple matrix grammars of finite index. Information Processing Letters, 7(2):100–102, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paun:1979:SLA

[Pău79] Gheorghe Păun. On Szilard’s languages associated to a matrix grammar. Information Processing Letters, 8(2):104–105, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pankov:1979:SEE

[PD79] P. S. Pankov and Si. L. Dolmatov. Substantiable evaluations by electronic computers and their application to one problem in combinatorial geometry. Information Processing Letters, 8(4):202–203, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peltola:1978:IMS

[PE78] Eero Peltola and Hannu Erkiö. Insertion merge sorting. Information Processing Letters, 7(2):92–99, February 28, 1978. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pearl:1975:CIC

[Pea75] Judea Pearl. On the complexity of inexact computations. Information Processing Letters, 4(3):77–81, December ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Penttonen:1975:EGG

[Pen75] Martti Penttonen. ET0L-grammars and N-grammars. Information Processing Letters, 4(1):11–13, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peterson:1980:NCP

[Pet80a] J. L. Peterson. A note on colored Petri nets. Information Processing Letters, 11(1):40–43, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pettorossi:1980:DAC

[Pet80b] Alberto Pettorossi. Derivation of an O(k^{2} log n) algorithm for computing order-k Fibonacci numbers from the O(k^{3} log n) matrix multiplication method. Information Processing Letters, 11(4–5):172–179, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pequeno:1979:ADT

[PL79] T. H. Pequeno and C. J. Lucena.

An approach for data type specification and its use in program verification. Information Processing Letters, 8(2):98–103, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pleszczynski:1975:GP

[Ple75] Stefan Pleszczyński. On the generation of permutations. Information Processing Letters, 3(6):180– 183, July ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Plesnik:1979:NCH

[Ple79] J. Plesník. The NP-completeness of the Hamiltonian cycle problem in planar digraphs with degree bound two. Information Processing Letters, 8(4):199–201, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Preparata:1975:TRE

[PM75] F. P. Preparata and D. E. Muller. The time required to evaluate division-free arithmetic expressions. Information Processing Letters, 3 (5):144–146, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Partsch:1976:FRR

[PP76] H. Partsch and P. Pepper. A family of rules for recursion removal. Information Processing Letters, 5(6):174– 177, December ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paredaens:1980:GLA

[PP80] J. Paredaens and F. Ponsaert. Grant levels in an authorization mechanism. Information Processing Letters, 11(4–5):152–155, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Perl:1977:UCI

[PR77] Yehoshua Perl and Edward M. Reingold. Understanding the complexity of interpolation search. Information Processing Letters, 6(6):219– 222, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pritchard:1976:PRM

[Pri76] P. A. Pritchard. A proof rule for multiple coroutine systems. Information Processing Letters, 4(6):141– 143, March ??, 1976. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Prini:1979:SIS

[Pri79] Gianfranco Prini. Stack implementation of shallow binding in languages with mixed scoping. Information Processing Letters, 9(3): 143–154, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Probert:1976:CNC

[Pro76] R. L. Probert. Commutativity, noncommutativity, and bilinearity. Information Processing Letters, 5(2): 46–49, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Preparata:1978:IPP

[PS78] F. P. Preparata and D. V. Sarwate. An improved parallel processor bound in fast matrix inversion. Information Processing Letters, 7(3):148–150, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petreschi:1980:SAS

[PS80] Rossella Petreschi and Bruno Simeone. A switching algorithm for the solution of quadratic Boolean equations. Information Processing Letters, 11(4–5):193–198, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [PS81].

Petreschi:1981:ESA

[PS81] R. Petreschi and B. Simeone. Erratum: “A switching algorithm for the solution of quadratic Boolean equations”. Information Processing Letters, 12(2):109, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [PS80].

Preparata:1980:AOV

[PV80] Franco P. Preparata and Jean E. Vuillemin. Area-time optimal VLSI networks for multiplying matrices. Information Processing Letters, 11 (2):77–80, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramshaw:1977:BCN

[Ram77] Lyle Ramshaw. Binomial coefficients with non-integral lower index. Information Processing Letters, 6(6):223–226, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ridjanovic:1982:DDD

[RB82] Dzenan Ridjanovic and Michael L. Brodie. Defining database dynamics with attribute grammars. Information Processing Letters, 14(3): 132–138, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reddy:1972:XHS

[RBE^{+}72] R. Reddy, W. Broadley, L. Erman, R. Johnsson, J. Newcomer, G. Robertson, and J. Wright. Xcribl — a hardcopy scan line graphics system for document generation. Information Processing Letters, 1(6):246– 251, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rick:1979:IBA

[RE79] C. C. Rick and D. J. Evans. An improved bisection algorithm. Information Processing Letters, 8(2):112– 113, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reddi:1977:ASC

[Red77] S. S. Reddi. Alternate solutions to the cigarette smokers’ problem without conditionals. Information Processing Letters, 6(1):22–24, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reiser:1978:LSA

[Rei78] A. Reiser. A linear selection algorithm for sets of elements with weights. Information Processing Letters, 7(3):159–162, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reising:1979:NRF

[Rei79a] Wolfgang Reising. A note on the representation of finite tree automata. Information Processing Letters, 8(5):239–240, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reiss:1979:RS

[Rei79b] Steven P. Reiss. Rational search. Information Processing Letters, 8(2): 89–90, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rem:1975:PES

[Rem75] M. Rem. On the programming of elastic stores. Information Processing Letters, 3(6):184–187, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reusch:1976:GLA

[Reu76] Peter J. A. Reusch. Generalized lattices applicable in retrieval models. Information Processing Letters, 4(6):144–148, March ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roschke:1973:AOC

[RF73] S. I. Roschke and A. L. Furtado. An algorithm for obtaining the chromatic number and an optimal coloring of a graph. Information Processing Letters, 2(2):34–38, June ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reiser:1975:EED

[RK75a] John F. Reiser and Donald E. Knuth. Erratum: “Evading the Drift in Floating-Point Addition”. Information Processing Letters, 3 (5):164, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [RK75b].

Reiser:1975:EDF

[RK75b] John F. Reiser and Donald E. Knuth. Evading the drift in floatingpoint addition. Information Processing Letters, 3(3):84–87, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [RK75a].

Rosenfeld:1973:PSA

[RM73] Azriel Rosenfeld and David L. Milgram. Parallel/sequential array automata. Information Processing Letters, 2(2):43–46, June ??, 1973. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1973:IAT

[Rob73] J. M. Robson. An improved algorithm for traversing binary trees without auxiliary stack. Information Processing Letters, 2(1):12–14, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1975:SSI

[Rob75] J. M. Robson. A simple solution to the interleaved memory bandwidth problem. Information Processing Letters, 3(4):125–126, March ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1979:ECP

[Rob79] J. M. Robson. The emptiness of complement problem for semi extended regular expressions requires c^{n} space. Information Processing Letters, 9(5):220–222, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1980:SAN

[Rob80] J. M. Robson. Storage allocation is NP-hard. Information Processing Letters, 11(3):119–125, November 18, 1980. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohlf:1978:PMS

[Roh78] F. James Rohlf. A probabilistic minimum spanning tree algorithm. Information Processing Letters, 7(1): 44–48, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohrich:1982:HQW

[Röh82] Johannes Röhrich. A hybrid of Quicksort with O(n log n) worst case complexity. Information Processing Letters, 14(3):119–123, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Romani:1980:SPP

[Rom80] Francesco Romani. Shortest-path problem is not harder than matrix multiplication. Information Processing Letters, 11(3):134–136, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also note [Mor81].

Rosenberg:1979:NPE

[Ros79] Arnold L. Rosenberg. A note on paths embedded in trees. Information Processing Letters, 8(5):272– 273, June 11, 1979. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rotem:1975:CBB

[Rot75] D. Rotem. On a correspondence between binary trees and a certain type of permutation. Information Processing Letters, 4(3):58–61, December ??, 1975. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Roussopoulos:1973:ADG

[Rou73] Nicholas D. Roussopoulos. A max(m, n) algorithm for determining the graph H from its line graph G. Information Processing Letters, 2 (4):108–112, October ??, 1973. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rowicki:1975:NOS

[Row75] Andrzej Rowicki. A note on optimal scheduling for two-processor systems. Information Processing Letters, 4(2):27–30, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rowicki:1977:NOP

[Row77] A. Rowicki. A note on optimal preemptive scheduling for twoprocessor systems. Information Processing Letters, 6(1):25–28, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rozenberg:1972:DPU

[Roz72a] G. Rozenberg. Direct proofs of the undecidability of the equivalence problem for sentential forms of linear context-free grammars and the equivalence problem for 0L systems. Information Processing Letters, 1(6):233–235, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rozenberg:1972:EPD

[Roz72b] G. Rozenberg. The equivalence problem for deterministic T0Lsystems is undecidable. Information Processing Letters, 1(5):201– 204, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Roz72c].

Rozenberg:1972:EEP

[Roz72c] G. Rozenberg. Erratum: “The equivalence problem for deterministic T0L-systems is undecidable”. Information Processing Letters, 1 (6):252, December ??, 1972. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Roz72b].

Rozenberg:1976:MES

[Roz76a] G. Rozenberg. More on ET0L systems versus random context grammars. Information Processing Letters, 5(4):102–106, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rozenberg:1976:SIG

[Roz76b] G. Rozenberg. On slicing of K-iteration grammars. Information Processing Letters, 4(5):127– 131, February ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raiha:1977:OAS

[RS77] K.-J. Räihä and M. Saarinen. An optimization of the alternating semantic evaluator. Information Processing Letters, 6(3):97–100, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ruohonen:1975:TRC

[Ruo75] K. Ruohonen. Three results of comparison between L languages with and without interaction. Information Processing Letters, 4(1):7–10, September ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ruohonen:1978:NLE

[Ruo78] K. Ruohonen. A note on language equations involving morphisms. Information Processing Letters, 7(5): 209–212, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ruohonen:1979:DFD

[Ruo79] Keijo Ruohonen. The decidability of the F0L-D0L equivalence problem. Information Processing Letters, 8(5):257–260, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajaraman:1977:FFB

[RV77] V. Rajaraman and Om Vikas.

A first-in-first-out buffered cyclic memory. Information Processing Letters, 6(2):63–68, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rivest:1979:LBN

[RV79] Ronald L. Rivest and Jean-Paul Van de Wiele. An Ω((n/ lg n)^{1/2}) lower bound on the number of additions necessary to compute 0-1 polynomials over the ring of integer polynomials. Information Processing Letters, 8(4):178–180, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rozenberg:1976:NIG

[RW76] G. Rozenberg and D. Wood. A note on K-iteration grammars. Information Processing Letters, 4(6): 162–164, March ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raghavan:1977:NMS

[RY77] V. V. Raghavan and C. T. Yu. A note on a multidimensional searching problem. Information Processing Letters, 6(4):133–135, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salton:1971:PII

[Sal71] G. Salton. The performance of interactive information retrieval. Information Processing Letters, 1(2):35– 41, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salton:1973:EML

[Sal73] G. Salton. Experiments in multilingual information retrieval. Information Processing Letters, 2(1): 6–11, March ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Samadi:1976:TSM

[Sam76] Behrokh Samadi. B-trees in a system with multiple users. Information Processing Letters, 5(4):107– 112, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Samet:1978:CFA

[Sam78] H. Samet. A canonical form algorithm for proving equivalence of conditional forms. Information Processing Letters, 7(2):103–106, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Santoro:1976:FTS

[San76] Nicola Santoro. Full table search by polynomial functions. Information Processing Letters, 5(3):72–74, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sengupta:1979:ICP

[SBS79] A. Sengupta, S. Bandyopadhyay, and P. K. Srimani. On identification of CR property in file organisation. Information Processing Letters, 9(2):93–96, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Scheiding:1971:REM

[Sch71] H. Scheiding. Representation and equality of modes. Information Processing Letters, 1(2):61–65, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schachtel:1978:NAM

[Sch78] G. Schachtel. A noncommutative algorithm for multiplying 5 × 5 matrices using 103 multiplications. Information Processing Letters, 7(4): 180–182, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwartz:1979:AAP

[Sch79] R. L. Schwartz. Aliasing among pointers in EUCLID. Information Processing Letters, 9(2):76–79, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schmitt:1982:CPF

[Sch82] Alfred Schmitt. On the computational power of the floor function. Information Processing Letters, 14 (1):1–3, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sekey:1975:GFE

[Sek75] A. A. Sekey. A generating function for entropy. Information Processing Letters, 4(2):51–??, November ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Selkow:1977:TTE

[Sel77] S. M. Selkow. The tree-to-tree editing problem. Information Processing Letters, 6(6):184–186, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Selkow:1978:NBC

[Sel78] S. M. Selkow. New bounds for the clique number of a graph. Information Processing Letters, 7(4): 173–174, June ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sethi:1973:NIP

[Set73] Ravi Sethi. A note on implementing parallel assignment instructions. Information Processing Letters, 2(4): 91–95, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Silva-Filho:1979:ACA

[SF79] Y. V. Silva-Filho. Average case analysis of region search in balanced k-d trees. Information Processing Letters, 8(5):219–223, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sassa:1976:HMF

[SG76] M. Sassa and E. Goto. A hashing method for fast set operations. Information Processing Letters, 5(2): 31–34, June ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sassa:1977:VTV

[SG77] Masataka Sassa and Eiichi Goto.

“V-tape”, a virtual memory oriented data type, and its resource requirements. Information Processing Letters, 6(2):50–55, April ??, 1977. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simovici:1979:EIF

[SG79] Dan A. Simovici and Gh. Grigoraş. Even initial feedback vertex set problem is NP-complete. Information Processing Letters, 8(2): 64–66, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shamir:1979:FNA

[Sha79] Adi Shamir. Factoring numbers in O(log n) arithmetic steps. Information Processing Letters, 8(1):28–31, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shiloach:1979:EDB

[Shi79a] Yossi Shiloach. Edge-disjoint branching in directed multigraphs. Information Processing Letters, 8 (1):24–27, January 2, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shiloach:1979:FEC

[Shi79b] Yossi Shiloach. A fast equivalencechecking algorithm for circular lists. Information Processing Letters, 8 (5):236–238, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shiloach:1979:SLO

[Shi79c] Yossi Shiloach. Strong linear orderings of a directed network. Information Processing Letters, 8(3): 146–148, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shortt:1978:IPC

[Sho78] Joseph Shortt. An iterative program to calculate Fibonacci numbers in O(log n) arithmetic operations. Information Processing Letters, 7(6): 299–303, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shyamasundar:1976:NLP

[Shy76] R. K. Shyamasundar. A note on linear precedence functions. Information Processing Letters, 5(3):81–??, August ??, 1976. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Siklossy:1972:FRO

[Sik72] L. Siklóssy. Fast and read-only algorithms for traversing trees without an auxiliary stack. Information Processing Letters, 1(4):149–152, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schnorr:1977:UHS

[SK77] C. P. Schnorr and H. Klupp. A universally hard set of formulae with respect to non-deterministic Turing acceptors. Information Processing Letters, 6(2):35–37, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slisenko:1982:CFG

[Sli82] A. O. Slisenko. Context-free grammars as a tool for describing polynomial-time subclasses of hard problems. Information Processing Letters, 14(2):52–56, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slowinski:1979:CMP

[Sło79] R. Słowiński. Cost-minimal preemptive scheduling of independent jobs with release and due dates on open shop under resource constraints. Information Processing Letters, 9(5): 233–237, December 16, 1979. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sokolowski:1978:MPP

[Sok78] S. Sokolowski. A method for proving programming languages non context-free. Information Processing Letters, 7(3):151–153, April 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Solntseff:1972:CEP

[Sol72] N. Solntseff. A classification of extensible programming languages. Information Processing Letters, 1(3): 91–96, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Solntseff:1974:NDD

[Sol74] N. Solntseff. On a notational device for the description of pointer-free operations on structured data. Information Processing Letters, 2(6): 158–159, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Soundararajan:1979:APT

[Sou79] N. Soundararajan. Axiomatic proofs of total correctness of programs. Information Processing Letters, 8(5): 274–277, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srinivasan:1979:NRD

[SR79] B. Srinivasan and V. Rajaraman. On the normalization of relational databases. Information Processing Letters, 9(2):89–92, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sridharan:1974:CGV

[Sri74] N. S. Sridharan. Computer generation of vertex graphs. Information Processing Letters, 3(2):57– 63, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Sri75].

Sridharan:1975:ECG

[Sri75] N. S. Sridharan. Erratum: “Computer Generation of Vertex Graphs”. Information Processing Letters, 3 (5):164, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Sri74].

Sethi:1978:ECC

[SRK78] A. S. Sethi, V. Rajaraman, and P. S.

Kenjale. An error-correcting coding scheme for alphanumeric data. Information Processing Letters, 7(2): 72–77, February 28, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Steckelberg:1977:RBA

[SS77] J. M. Steckelberg and S. C. Seth.

On a relation between algebraic programs and Turing machines. Information Processing Letters, 6(6):180– 183, December ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stoyan:1979:MMS

[SS79] Yu. G. Stoyan and V. Z. Socolovsky.

The minimization method for some permutation functionals. Information Processing Letters, 8(2):110– 111, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shaw:1977:SGA

[ST77] M. Shaw and J. F. Traub. Selection of good algorithms from a family of algorithms for polynomial derivative evaluation. Information Processing Letters, 6(5):141–145, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

SangAjang:1979:EAD

[ST79] Gerard J. M. Sang Ajang and Frank Teer. An efficient algorithm for detection of combined occurrences. Information Processing Letters, 8(3): 137–140, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stacey:1971:RVM

[Sta71] G. M. Stacey. The role of virtual memory in the handling of application files. Information Processing Letters, 1(1):1–3, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stevens:1975:ECA

[Ste75] A. Stevens. An elementary computer algorithm for the calculation of the coefficient of inbreeding. Information Processing Letters, 3(5): 153–163, May ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sudborough:1977:NWO

[Sud77] I. H. Sudborough. A note on weak operator precedence grammars. Information Processing Letters, 6(6): 213–218, December ??, 1977. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sudborough:1978:NWO

[Sud78] I. H. Sudborough. A note on weak operator precedence grammars. Information Processing Letters, 7(5):

213–218, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Solntseff:1972:EEC

[SY72] N. Solntseff and A. Yezerski. ECT:

an extensible-contractable translator system. Information Processing Letters, 1(3):97–99, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szwarcfiter:1979:SDR

[Szw79] J. L. Szwarcfiter. Systems of distinct representatives for k families of sets. Information Processing Letters, 9(5):195–196, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takaoka:1974:NAC

[Tak74] T. Takaoka. A note on the ambiguity of context-free grammars. Information Processing Letters, 3(2): 35–36, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Talbert:1977:FSS

[Tal77] S. H. Talbert. On the formal specification of the semantics of processed information. Information Processing Letters, 6(4):116–119, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1972:DWG

[Tar72] R. E. Tarjan. Determining whether a groupoid is a group. Information Processing Letters, 1(3):120– 124, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1974:GAE

[Tar74a] R. E. Tarjan. A good algorithm for edge-disjoint branching. Information Processing Letters, 3(2):51– 53, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1974:NFB

[Tar74b] R. E. Tarjan. A note on finding the bridges of a graph. Information Processing Letters, 2(6):160–161, April ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1974:NAF

[Tar74c] Robert Endre Tarjan. A new algorithm for finding weak components. Information Processing Letters, 3 (1):13–15, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarhio:1982:LPS

[Tar82a] Jorma Tarhio. LR parsing of some ambiguous grammars. Information Processing Letters, 14(3):101–103, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1982:HCA

[Tar82b] R. Endre Tarjan. A hierarchical clustering algorithm using strong components. Information Processing Letters, 14(1):26–29, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1982:SAM

[Tar82c] Robert Endre Tarjan. Sensitivity analysis of minimum spanning trees and shortest path trees. Information Processing Letters, 14(1): 30–33, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [Tar86].

Tarjan:1986:SAM

[Tar86] R. E. Tarjan. Corrigendum: “Sensitivity analysis of minimum spanning trees and shortest path trees” [Inform. Process. Lett. 14(1), 27 March 1982, pp. 30–33]. Information Processing Letters, 23(4):219, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Tar82c].

tenHoopen:1980:CRR

[ten80] J. ten Hoopen. Consecutive retrieval with redundancy: an optimal linear and an optimal cyclic arrangement and their storage space requirements. Information Processing Letters, 11(4–5):211–217, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

terBekke:1980:CD

[ter80] J. H. ter Bekke. Convertibility in databases. Information Processing Letters, 11(4–5):168–171, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teskey:1979:DRU

[Tes79] F. N. Teskey. Document retrieval using associative processors. Information Processing Letters, 9(2):64–67, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teuhola:1978:CMC

[Teu78] Jukka Teuhola. A compression method for clustered bit-vectors. Information Processing Letters, 7(6): 308–311, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Terashima:1978:GOC

[TG78] Motoaki Terashima and Eiichi Goto. Genetic order and compactifying garbage collectors. Information Processing Letters, 7(1):27–32, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Toussaint:1972:FEQ

[Tou72] G. T. Toussaint. Feature evaluation with quadratic mutual information. Information Processing Letters, 1(4):153–156, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trivedi:1976:SA

[Tri76] K. S. Trivedi. On a semaphore anomaly. Information Processing Letters, 5(3):88–89, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsao:1981:NIB

[Tsa81] Nai Kuan Tsao. The numerical instability of Bini’s algorithm. Information Processing Letters, 12(1): 17–19, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See reply [Bin82].

Tsichritzis:1971:NCS

[Tsi71] D. Tsichritzis. A note on comparison of subrecursive hierarchies. Information Processing Letters, 1(2):42–44, July ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsichritzis:1972:POS

[Tsi72] D. Tsichritzis. Protection in operating systems. Information Processing Letters, 1(4):127–131, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsur:1978:AQN

[Tsu78] Shalom Tsur. Analysis of queuing networks in which processes exhibit locality-transition behaviour. Information Processing Letters, 7(1): 20–23, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turner:1977:EDR

[Tur77] D. A. Turner. Error diagnosis and recovery in one pass compilers. Information Processing Letters, 6(4): 113–115, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tang:1973:MBB

[TW73] D. T. Tang and C. K. Wong. A modified branch-and-bound strategy. Information Processing Letters, 2(3): 65–69, August ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzafestas:1972:DPM

[Tza72a] S. G. Tzafestas. Design parameters for a multiserver computer processing buffering system with feedback. Information Processing Letters, 1(6):236–243, December ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzafestas:1972:IOM

[Tza72b] S. G. Tzafestas. Input-output modeling and identification of linear automata. Information Processing Letters, 1(3):69–75 (or 69–70??), February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ukkonen:1979:NSC

[Ukk79] Esko Ukkonen. The nonexistence of some covering context-free grammars. Information Processing Letters, 8(4):187–192, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Umeo:1982:DOW

[UMS82] Hiroshi Umeo, Kenichi Morita, and Kazuhiro Sugata. Deterministic one-way simulation of two-way realtime cellular automata and its related problems. Information Processing Letters, 14(4):158–161, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Upfal:1982:FCP

[Upf82] Eliezer Upfal. Formal correctness proofs of a nondeterministic program. Information Processing Letters, 14(2):86–92, April 20, 1982.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Urbanek:1980:ACE

[Urb80] Friedrich J. Urbanek. An O(log n)

algorithm for computing the n th element of the solution of a difference equation. Information Processing Letters, 11(2):66–67, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valiant:1976:RCC

[Val76] L. G. Valiant. Relative complexity of checking and evaluating. Information Processing Letters, 5(1):20– 23, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valiant:1980:CMP

[Val80] L. G. Valiant. Computing multivariate polynomials in parallel. Information Processing Letters, 11(1):44–45, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanLeeuwen:1975:MQE

[Van75] Jan Van Leeuwen. The membership question for ET0L-languages is polynomially complete. Information Processing Letters, 3(5):138– 143, May ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanEmdeBoas:1977:POF

[van77] P. van Emde Boas. Preserving order in a forest in less than logarithmic time and linear space. Information Processing Letters, 6(3):80–82, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanderNat:1979:BMP

[Van79a] M. Van Der Nat. Binary merging by partitioning. Information Processing Letters, 8(2):72–75, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeeuwen:1979:CSD

[van79b] Jan van Leeuwen. On compromising statistical data-bases with a few known elements. Information Processing Letters, 8(3):149–153, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanEmdeBoas:1980:LBC

[van80] Peter van Emde Boas. On the Ω(n log n) lower bound for convex hull and maximal vector determination. Information Processing Letters, 10(3):132–136, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [Avi80].

Varney:1972:PPU

[Var72] R. C. Varney. Priority processes used for scheduling within a tree structured operating system. Information Processing Letters, 1(5):187– 190, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Varhegyi:1973:NDE

[Vár73] G. Várhegyi. Numerical differentiation of experimental data. Information Processing Letters, 2(1):24–25, March ??, 1973. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Veloso:1979:CRP

[Vel79] P. A. S. Veloso. Characterizing the regular prefix codes and right power-bounded languages. Information Processing Letters, 9(1):5–7, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vikas:1979:API

[Vik79] Om Vikas. Analysis of a periodically inspected buffer. Information Processing Letters, 8(3):124–130, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanderKnijff:1979:CCB

[VL79] D. Van der Knijff and J.-L. Lassez.

A clarification of the comparison between some measures of software science. Information Processing Letters, 8(5):241–243, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeeuwen:1982:ERR

[vN82] Jan van Leeuwen and Maurice Nivat. Efficient recognition of rational relations. Information Processing Letters, 14(1):34–38, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vonSolms:1975:TLT

[Von75] S. H. Von Solms. On T0L languages over terminals. Information Processing Letters, 3(3):69–70, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeeuwen:1974:IBD

[vS74] J. van Leeuwen and C. H. Smith. An improved bound for detecting looping configurations in deterministic DPA’s. Information Processing Letters, 3(1):22–24, July ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanDam:1972:SDP

[vT72] A. van Dam and F. Wm. Tompa. Software data paging and segmentation for complex systems. Information Processing Letters, 1(3):80– 86, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vantilborgh:1972:EDS

[Vv72] H. Vantilborgh and A. van Lamsweerde. On an extension of Dijkstra’s semaphore primitives. Information Processing Letters, 1(5): 181–186, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeeuwen:1980:ELS

[vW80a] J. van Leeuwen and D. Wood. Errata: “List scheduling bounds for UET systems with resources” [Inform. Process. Lett. 10 (1980), no. 1, 28–31; MR 81a:68042]. Information Processing Letters, 11(1):57, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [vW80b].

vanLeeuwen:1980:DDS

[vW80b] Jan van Leeuwen and Derick Wood. Dynamization of decomposable searching problems. Information Processing Letters, 10(2):51–56, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [vW80a].

Wang:1976:ERM

[Wan76a] Patrick Shen-Pei Wang. Erratum:

“Recursiveness of Monotonic Array Grammars and a Hierarchy of Array Languages”. Information Processing Letters, 5(3):90, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Wan76b].

Wang:1976:RMA

[Wan76b] Patrick Shen-Pei Wang. Recursiveness of monotonic array grammars and a hierarchy of array languages. Information Processing Letters, 5 (1):24–26, May ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Wan76a].

Warren:1973:MCS

[War73] H. S. Warren, Jr. Minimal comparison sorting by choosing most efficient comparisons. Information Processing Letters, 2(5):129–130, December ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1979:TCG

[WB79] M. H. Williams and A. R. Bulmer.

A transportable code generator generator system. Information Processing Letters, 9(3):122–125, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegbreit:1975:RCT

[Weg75] B. Wegbreit. Retrieval from context trees. Information Processing Letters, 3(4):119–120, March ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weihrauch:1975:PSP

[Wei75] Klaus Weihrauch. Program schemata with polynomial bounded counters. Information Processing Letters, 3(3):91–96, January ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wessner:1976:OAS

[Wes76] Russell L. Wessner. Optimal alphabetic search trees with restricted maximal height. Information Processing Letters, 4(4):90–94, January ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Whale:1972:CVB

[Wha72] J. F. Whale. The critical value of the basic parameter of a nonlinear differential equation. Information Processing Letters, 1(3):87–90, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1977:COP

[Wil77] M. H. Williams. Complete operator precedence conditions. Information Processing Letters, 6(2):60–62, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1980:BSB

[Wil80a] M. H. Williams. Batch sizes for the batching method of colouring planar maps. Information Processing Letters, 11(4–5):186–189, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1980:CMC

[Wil80b] M. H. Williams. Cubic map configurations. Information Processing Letters, 11(4–5):180–185, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Winkowski:1977:ACB

[Win77] Jozef Winkowski. An algebraic characterization of the behaviour of non-sequential systems. Information Processing Letters, 6(4):105– 109, August ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also errata [Win78].

Winkowski:1978:EAC

[Win78] J. Winkowski. Erratum: “An algebraic characterization of the behaviour of non-sequential systems”. Information Processing Letters, 7 (1):63, January 12, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Win77].

Watjen:1982:AVE

[WS82] Dietmar Wätjen and Werner Struckmann. An algorithm for verifying equations of morphisms in a category. Information Processing Letters, 14(3):104–108, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wurgens:1977:CER

[Wür77] H. Würgens. Comments on “Error resynchronisation in producer consumer systems”. Information Processing Letters, 6(3):87–90, April ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1975:AFM

[Yao75a] A. C. Yao. An O(|E|) log log |V | algorithm for finding minimum spanning trees. Information Processing Letters, 4(1):21–23, September ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1975:VAF

[Yao75b] Andrew C. Yao. An O(|E|) log log |V | algorithm for finding minimum spanning trees. Information Processing Letters, 4(??):21–23, ?? ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1979:NCK

[Yao79a] Andrew C. Yao. A note on a conjecture of Kam and Ullman concerning statistical databases. Information Processing Letters, 9(1):48–50, July 20, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1979:GIN

[Yao79b] F. Frances Yao. Graph 2isomorphism is NP-complete. Information Processing Letters, 9(2): 68–72, August 17, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1980:NAE

[Yao80] Andrew C. Yao. A note on the analysis of extendible hashing. Information Processing Letters, 11(2):84–86, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yehudai:1979:NPL

[Yeh79] Amiram Yehudai. A note on the pumping lemma for regular languages. Information Processing Letters, 9(3):135–136, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:1976:CFS

[YJ76] C. T. Yu and D. T. Johnson. On the complexity of finding the set of candidate keys for a given set of functional dependencies. Information Processing Letters, 5(4):100– 101, October ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:1979:NAN

[YL79] C. C. Yang and D. T. Lee. A note on the all nearest-neighbor problem for convex polygons. Information Processing Letters, 8(4):193–194, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LP78, FK79].

Yuval:1973:PCT

[Yuv73] G. Yuval. On a problem connected with topological sorting. Information Processing Letters, 2(4):103– 104, October ??, 1973. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuval:1975:FNN

[Yuv75] G. Yuval. Finding near neighbours in K-dimensional space. Information Processing Letters, 3(4):113– 114, March ??, 1975. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuval:1976:AFA

[Yuv76a] G. Yuval. An algorithm for finding all shortest paths using N^{2.81} infinite-precision multiplications. Information Processing Letters, 4(6):155–156, March ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuval:1976:FNN

[Yuv76b] G. Yuval. Finding nearest neighbours. Information Processing Letters, 5(3):63–65, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuval:1977:TE

[Yuv77] G. Yuval. Theil’s estimator. Information Processing Letters, 6(1): 6–7, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuval:1978:SPS

[Yuv78] G. Yuval. A simple proof of Strassen’s result. Information Processing Letters, 7(6):285–286, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [Yuv79].

Yuval:1979:CSP

[Yuv79] G. Yuval. Corrigendum: “A simple proof of Strassen’s result” (Inform. Process. Lett. 7 (1978), no. 6, 285–286). Information Processing Letters, 8(5):278, June 11, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Yuv78].

Yokomori:1982:TRN

[YWL82] Takashi Yokomori, Derick Wood, and Klaus-Jörn Lange. A threerestricted normal form theorem for ETOL languages. Information Processing Letters, 14(3):97–100, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [YWL85].

Yokomori:1985:ETR

[YWL85] Takashi Yokomori, Derick Wood, and Klaus-Jörn Lange. Erratum: “A three-restricted normal form theorem for ETOL languages” [Inform. Process. Lett. 14 (1982), no. 3, 97– 100; MR 83g:68115]. Information Processing Letters, 21(1):53, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [YWL82].

Zaks:1982:GRA

[Zak82] Shmuel Zaks. Generation and ranking of K-ary trees. Information Processing Letters, 14(1):44–48, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zave:1975:FCG

[Zav75] D. A. Zave. A fast compacting garbage collector. Information Processing Letters, 3(6):167–169, July ??, 1975. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zave:1976:SEI

[Zav76] D. A. Zave. A series expansion involving the harmonic numbers. Information Processing Letters, 5(3): 75–77, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zissos:1971:PSR

[ZD71] D. Zissos and F. G. Duncan. Programmed simulation of race hazards in sequential circuits. Information Processing Letters, 1(1):11– 13, February ??, 1971. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zelkowitz:1974:SOS

[Zel74] M. V. Zelkowitz. Structured operating system organization. Information Processing Letters, 3(2):39– 42, November ??, 1974. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zimmer:1972:SP

[Zim72] R. Zimmer. Soft precedence. Information Processing Letters, 1(3):108– 110, February ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zuczek:1977:USP

[Zuc77] R. Zuczek. The universal space for parallel computation. Information Processing Letters, 6(2):42–45, April ??, 1977. CODEN IFPLAT. ISSN REFERENCES 84

0020-0190 (print), 1872-6119 (electronic).